### v1

# Mitigating Polarization in Recommender Systems via Network-aware Feedback Optimization

Abstract
We consider a recommender system that takes into account the interaction between recommendations and the evolution of user interests. Users opinions are influenced by both social interactions and recommended content. We leverage online feedback optimization to design a recommender system that trades-off between maximizing engagement and minimizing polarization. The recommender system is agnostic about users’ opinion, clicking behavior, and social interactions, and solely relies on clicks. We establish optimality and closed-loop stability of the resulting feedback interconnection between the social platform and the recommender system. We numerically validate our algorithm when the user population follows an extended Friedkin–Johnsen model. We observe that network-aware recommendations significantly reduce polarization without compromising user engagement.
{IEEEkeywords}
Recommender systems, networked control systems, opinion dynamics, non-convex optimization.
1 Introduction
Problem Description
Online social platforms use recommender systems to provide users with tailored content to maximize engagement over the platform. To do so, recommender systems exploit methods, such as content-based filtering [1] and collaborative filtering [2], that combine information personalization, popularity and similarity of interests with other users. An important drawback of content personalization is related to the amplification phenomenon [3] associated with the feedback loop that arises between recommended content and user preferences. Amplification leads to misinformation spread [4], filter bubbles [5], bias exacerbation [6] and ultimately polarization [7, 8]. Driven by these considerations we propose a recommender system that takes into account the interplay between recommendations and user interests. Our goal is to understand i) the impact of recommendations on user opinions and ii) how should recommendation system depart from engagement maximization to mitigate polarization. By leveraging on online feedback optimization (OFO) [9] and building upon [10, 11], we design a network-aware, multiple objective, recommender system as a feedback controller, illustrated in Figure 1.
[…]
Figure 1: Illustration of the closed-loop between the opinion dynamics and the recommender system.
Related Work
Recommendation systems. Most of the works on recommender system are static (e.g. matrix factorization) and neglect the dynamic nature of the feedback interaction between the user and the online platforms. Our work develops along a recent systems-theory perspective on recommender systems [10] where the feedback interaction between social platform and recommender systems is made explicit. In [10], the recommender system provides only extreme positions to the user based on its clicking history. They show that when confirmation bias [12], namely the tendency to engage with content that aligns with prior beliefs, is the only driving force to click on contents, the recommender system induces opinion polarization. The recommender design proposed in [13] alternates between exploitation and exploration by either providing news which more closely align with the users’ beliefs or exploring new content. They show that shifts in the opinions at the individual level do not necessarily reflect in the opinion distribution of the population. The work [14] models the recommender system as an artificial user in the network that provides news content at minimum distance from all users opinions to promote engagement. The authors study how recommender systems of this form affect opinion dynamics. In [11], they consider the interplay between one single user, the recommender and harmful content and seek recommendation policies that tradeoff between maximizing the click-through rate (CTR) and mitigate harm. Their finding suggest that to properly mitigate harm, it is necessary to account for user dynamics. Importantly, most of the works related to recommender systems largely focus on one single user [11, 13, 10] without explicitly considering the influence of social interactions.
Online feedback optimization. Conventional OFO algorithms employ first-order gradient methods [15] that require knowledge about real-time states, sensitivity of the input-output steady-state map and the gradients of the objective function with respect to the input and the states. Recent advancements [16, 17, 18, 19], however allow the employment of OFO with less information. In particular, in [16] the sensitivity matrix is estimated online using online state information. In [17, 18] an OFO scheme is proposed where supervised learning is employed for real-time states and gradient estimation, respectively. Finally, in [19], zeroth-order methods are used to directly estimate the composite gradient using cost function evaluations, thus not requiring sensitivity estimation. See Table 1 for an overview of the recent advancements in OFO.
Table 1: State-of-the-art OFO algorithms.
OFO Method Sensitivity States Cost function Gradients Convexity
[15]
[16]
[17]
[18, Alg. 2]
[19]
Our work
Contributions
Recommender systems provide users tailored content based on implicit feedback, such as clicks. We design an OFO-based recommender system algorithm that extracts information on state, sensitivity and objective function gradient from clicks and dynamically provides user-tailored recommendations.
The recommender system encompasses polarization mitigation and engagement maximization in the objective function. We show that by incorporating information about real-time opinions and clicking behaviors of users as well as the topology of the social network, the recommender is able to enhance engagement while mitigating polarization.
The recommendation algorithm is interpreted as a controller that incorporates opinion (state), topology (sensitivity), and clicking behavior (gradient) estimations. The system dynamics is assumed to be uniformly exponentially stable for constant inputs and to admit a continuously differentiable steady-state map. We provide theoretical closed-loop guarantees between the opinion dynamics and the recommender. Numerical results show that our network-aware recommender system compares favourably in terms of polarization with negligible engagement losses against isolated user scenarios, where the recommendations are solely provided based on personalization and network interactions are neglected.
Our contribution is threefold:
  1. 1.
    We propose a recommender system algorithm that takes into account users’ preferences and trades off between maximizing CTR and minimizing polarization. The recommender system is agnostic about real-time opinions, network interactions and clicking behaviour.
  2. 2.
    From a technical perspective, the OFO algorithm runs combining estimations of system state, input-output sensitivity, and objective function gradient on a non-convex optimization problem. As an independent technical contribution, we provide convergence guarantees towards a locally optimal solution, thereby improving the state of the art on feedback optimization.
  3. 3.
    We validate the proposed control scheme on synthetic data generated via an extension of the Friedkin–Johsnen (FJ) model [10] that also takes into account the recommended content influence over opinions. We show the importance of network awareness on the attenuation of opinion polarization.
Outline
In Section 2, we formulate the optimization problem for engagement maximization and polarization mitigation. In Section 3, we describe the recommender design using OFO. In Section 4, we provide numerical simulations to show our algorithm performances and illustrate the benefits of network-awareness. Section 5 concludes our work.
Notations and Preliminaries
We let the symbols $\mathbb{R}$($\mathbb{R}_{+}$), $\mathbb{N}_{0}$ denote the set of (positive) real numbers and non-negative integers, respectively. The set of integers $\{1,2,\ldots,n\}$ is denoted by $[n]$. For a vector $y\in\mathbb{R}^{n}$, we let the symbol $|y|$ denote the vector whose $i$-th entry, $|y|_{i}$, is the $i$-th component of $y$ in modulus, $|y_{i}|$. The symbol $\langle\cdot,\cdot\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ denotes the standard inner product on $\mathbb{R}^{n}$. The value of a signal $x$ at time instant $k$ is denoted by $x^{k}$. The symbol $1_{n},(0_{n})$, denotes the all ones (zeros) vector of size $n$. The symbols $I_{n}$ and $O_{n}$, denote the $n$-dimensional identity and zero matrix, respectively. Given a matrix $A\in\mathbb{R}^{m\times n}$, the matrix is positive (non-negative) if all its elements are greater (or equal) than zero and we denote it as $A>0$ ($A\geq 0$). We let $\mathrm{vec}(A)$ denote the vectorized version of $A$, i.e., if $A=[a_{1}\ a_{2}\ \ldots\ a_{n}]$, with $a_{j}\in\mathbb{R}^{m},j\in[n]$, then $\mathrm{vec}(A)=[a_{1}^{\top}a_{2}^{\top}\ldots a_{n}^{\top}]^{\top}$. For a matrix $A\in\mathbb{R}^{n\times n}$, we denote by $\mathrm{det}[A]$ its determinant; $A$ is row-stochastic if $A\geq 0$ and $A1_{n}=1_{n}$. The symbol $\mathcal{D}_{\mathcal{U}}\subseteq\mathbb{R}$ denotes the set of $n$-dimensional diagonal matrices with diagonal entries in $\mathcal{U}$. The symbol $\mathrm{diag}[x]$ denotes the diagonal matrix with $x_{i}$ on its $i$-th diagonal entry. Given $\mathcal{C}\subset\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, we let $\Pi_{\mathcal{C}}[x]$ denote the Euclidean projection of $x$ over $\mathcal{C}$. The normal distribution is denoted by $\mathcal{N}(\mu,\Sigma)$, with $\mu,\Sigma$ representing the mean and variance, and the uniform distribution over the interval $[a,b]$, with $a,b,\in\mathbb{R}$ is represented by $\mathcal{U}[a,b]$.
We now state some preliminary lemmas and definitions that are recurrently used in this manuscript.
Definition 1 (Minimal modulus of continuity[20])
Let $h:\mathcal{X}\to\mathbb{R}$ be a continuous function over a bounded set $\mathcal{X}$. The minimal modulus of continuity of $h$ on $\mathcal{X}$ is defined as
$\omega_{h}(\gamma):=\mathrm{sup}\Big{\{}|h(x)-h(y)|:x,y\in\mathcal{X},\lVert x-y\rVert_{\infty}\leq\gamma\Big{\}}.$
Definition 2 (Persistently exciting input [21])
The signal $u\in\mathbb{R}^{n}$ is said to be persistently exciting of order $n$ if there exists a time span $S>0$ such that for all $k>0$, the matrix formed by the columns $u^{k+i}$ for $i\in\{0,1,\ldots S\}$ has full rank, i.e. $\mathrm{Rank}[u^{k},u^{k+1},\ldots,u^{k+S}]=n$.
Definition 3 (Global $\beta$-smoothness [22])
A continuously differentiable map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is globally $\beta$-smooth if its gradient $\nabla\Phi$ is $\beta$-Lipschitz.
Lemma 1 (Properties of $\beta$-smooth functions [22])
Given a globally $\beta$-smooth map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$, then $\Phi(x_{1})-\Phi(x_{2})-\nabla\Phi^{\top}(x_{2})(x_{1}-x_{2})\leq\frac{1}{2}\beta\lVert x_{1}-x_{2}\rVert^{2}$. Moreover, if $\Phi$ is twice continuously-differentiable, then $\lVert\nabla^{2}\Phi\rVert\leq\beta$.
2 Problem setup
2.1 Social Network Modelling
We consider a recommender system for a social network consisting of $n$ users, indexed by $i\in[n]$ that makes sequential, personalized recommendations. User opinions are collected into a vector $x\in[-1,1]^{n}$, with $x_{i}$, $i\in[n]$, denoting the opinion of the $i$-th user. The users’ opinions evolve as
$x^{k+1}=f(x^{k},p^{k},d),$ (1)
[…]
The following assumption ensures that the dynamics (1) are well-posed and admit a steady-state mapping.
Assumption 1 (Well-posedness)
The following hold:
  1. (i)
[…]
  2. (ii)
[…]
  3. (iii)
    The map $h(p,d)$ is continuously-differentiable and $L-$Lipschitz [22] with respect to $p$. $\square$
This assumption can be easily adopted for a wide array of renowned opinion dynamics models, such as De-Groot model [23], FJ model [24], and polar opinion dynamics model [25], among others, for which the influence of a recommendation system can be, in principle, included. For an extensive list of such models, we refer the reader to the survey [26]. Traditionally, in fact, these models consider social interactions and prejudice as the only driving forces behind opinion formation. Nonetheless, recommendations on contentious topics provably influence user opinions [10], underscoring the need to include their influence in the opinion dynamics. Next, we illustrate an example of how the FJ model, one of the most widely used opinion dynamics models across different communities [27, 28, 29, 30, 31, 10, 11], can be extended to capture the effects of the recommender on networked users.
Example 1 (Extended Friedkin–Johnsen model)
The opinions evolve according to
$x^{k+1}=(I_{n}-\Gamma_{p}-\Gamma_{d})Ax^{k}+\Gamma_{p}p^{k}+\Gamma_{d}d,$ (2)
where $\Gamma_{d},\Gamma_{p}$ are positive diagonal matrices such that $\Gamma_{p}+\Gamma_{d}\preceq I_{n}$, describing the impact of $d$ and $p$ over the opinions, and $A$ is a row-stochastic adjacency matrix encoding the social interconnections of the users. The dynamics (2) are forward-invariant in $[-1,1]^{n}$, since the opinions are a convex combination of $x,p,d\in[-1,1]^{n}$. Further, for a constant position provided by the recommender system, the steady-state mapping is single-valued, affine (hence, continuously differentiable), and reads as
$h(p,d)=(I_{n}-(I_{n}-\Gamma_{p}-\Gamma_{d})A)^{-1}(\Gamma_{p}p+\Gamma_{d}d).$
Note that by taking $\Gamma_{p}=O_{n}$ and $d=x^{0}$ to be an initial bias, the dynamics (2) reduce to the standard FJ model [24]. $\triangle$
Engagement of the users over the platform is measured in terms of CTR, measured as number of clicks over number of provided news articles. As common practice in the context of implicit feedback recommendation systems [32], for a generic user $i\in[n]$ holding an opinion $x_{i}$, we model the probability of $i$ clicking on a recommendation with position $p_{i}$ as a random variable $c_{i}$ sampled from a Bernoulli distribution with unknown argument $g_{i}(p_{i},x_{i})$, i.e. $c_{i}\sim\mathcal{B}(g_{i}(p_{i},x_{i}))$. The argument $g_{i}(p_{i},x_{i})$ represents the clicking behaviour of user $i$ and depends on the user’s opinion $x_{i}$ as well as the position of the recommendation $p_{i}$. We will work under the following regularity assumptions on the clicking behaviour:
Assumption 2 (Clicking behaviour)
The clicking behaviour $g(p,x):[-1,1]^{n}\times[-1,1]^{n}\to[0,1]^{n}$ is $M_{x}$-Lipschitz with respect to $x$, and $L_{p}$- and $L_{x}$-smooth with respect to $p$ and $x$, respectively. $\square$
A concrete example of a clicking behaviour is provided.
Example 2 (Extremity confirmation bias [10])
A user $i\in[n]$, holding opinion $x_{i}$, affected by extremity confirmation bias clicks on a recommendation $p_{i}$ with probability
$c_{i}\sim\mathcal{B}\left(\frac{1}{2}+\frac{1}{2}x_{i}p_{i}\right),$
where the clicking behaviour $g_{i}(p_{i},x_{i})=\frac{1}{2}+\frac{1}{2}x_{i}p_{i}$ models confirmation bias towards extreme recommendations. In fact, if the opinion $x_{i}\approx\pm 1$ and the position $p_{i}\approx\pm 1$ ($\mp 1$), then the probability of clicking is almost $1$ (00). The click event becomes random as the user position approaches the neutral stance $p_{i}\approx 0$, highlighting diminished engagement for less polarized content or when recommendations directly counter the user’s stance on the issue. $\triangle$
2.2 Problem formulation
The goal of the recommender system is to provide recommendations that optimize a specific metric, denoted as $\varphi(p,x)$. Typical choices include maximizing CTR [10, 11], opinion control [33], and polarization mitigation [34]. Deviating from traditional recommendation systems, we consider a multi-objective cost function combining engagement maximization and polarization mitigation, given by
$\varphi(p,x)=\varphi^{\text{ctr}}(p,x)+\gamma\,\varphi^{\text{pol}}(x),$ (3)
where $\gamma$ is a non-negative trade-off parameter. The engagement-related term in the cost (3) is the CTR, formalized as
$\varphi^{\text{ctr}}(p,x)=-\sum_{i\in[n]}\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,],$
where $\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,]$ is the expectation of user $i$’s clicking, given their opinion $x_{i}$ and a recommendation with position $p_{i}$. The second term in (3), $\varphi^{\rm pol}$, is a smooth Lipschitz function that measures opinion polarization. For our algorithm we consider it to be $\varphi^{\text{pol}}(x):=\sum_{i\in[n]}s_{i}(x)$, where $s_{i}$ is a soft penalty function:
$s_{i}(x_{i})=\begin{cases}(x_{i}-\epsilon_{1})^{2}&x_{i}<\epsilon_{1}\\ 0&\epsilon_{1}\leq x_{i}\leq\epsilon_{2}\\ (\epsilon_{2}-x_{i})^{2}&x_{i}>\epsilon_{2}\end{cases}$ (4)
Therein, the parameters $\epsilon_{1}\leq\epsilon_{2}$ are used to control the degree of penalty towards extreme opinions. Specifically, a smaller positive (negative) choice for $\epsilon_{2}$ ($\epsilon_{1}$) indicates a higher penalty on extreme positive (negative) opinions. According to [35], the metric in (4) falls in the class of polarization spread, indicating how extreme the opinions are. Note that, for $\epsilon_{1}=\epsilon_{2}=0$, the term (4) reduces to $\varphi^{\text{pol}}(x)=\lVert x\rVert^{2}$, a classic measure for opinion polarization [28].
The recommender system aims at finding the recommendation that drives the user population with dynamics (1) to a steady-state minimizing the multi-objective cost (3):
$\underset{p,\,x}{\mathrm{minimize}}$ $\quad\varphi(p,x)$ (5a)
s.t. $\quad x=h(p,d)$ (5b)
$\quad p\in[-1,1]^{n}$ (5c)
This problem is uncertain and potentially non-convex, since the analytical form of the clicking behaviour is unknown and likely complex. Likewise, the opinion dynamics (1) are unknown. If the opinion dynamics in (1) and the clicking behavior $g_{i}$ were known, and an accurate prediction of external influences $d$ were available, then one could aim to solve (5) offline. However, in practice, none of these information is readily available, and the recommender system provides positions solely based on clicks $\{c_{i}^{k}\}$.
Remark 1
Equation (5b) translates into the users’ opinion stationarity before the recommender system action is taken. This is consistent with most of the works on recommender systems based on preference dynamics[11, 10, 36, 6, 37], where the action is taken based on the steady-state dynamics of the proposed model.
Our feedback control problem is formally stated as follows:
Problem 1
Design a feedback controller $p(c)$ so that (1) converges to a solution $(p^{*},x^{*})$ of the optimization problem (5), assuming that only clicks $c_{i}^{k}$ are available. The clicking behaviour $g_{i}(p_{i},x_{i})$, steady-state map $h(p,d)$, the opinion dynamics (1), external influence $d$ are unknown, and the real-time opinions $x$ are not measurable.
3 Recommender system design
We approach Problem 1 by designing a dynamic feedback controller inspired by the projected-gradient descent algorithm in [38, Section 3A]. The resulting recommender system dynamically generates recommendations as
$p^{k+1}=\Pi_{[-1,1]^{n}}\big{[}p^{k}-\eta\,\Phi(p^{k},x^{k+1})\big{]},\quad\forall k\in\mathbb{N}$ (6)
where $\eta$ is a tunable controller gain, and
$\Phi(p,x):=\nabla_{p}\varphi(p,x)+\nabla_{p}h(p,d)^{\top}\nabla_{x}\varphi(p,x)$ (7)
represents the gradient obtained by applying the chain-rule of differentiation to the cost $\varphi(p,x)$ in (5), with opinions at steady state, i.e., $x=h(p,d)$. In practice, evaluating the gradient (7) at each sampling instant $k$ requires access to:
  1. (i)
    Real-time users’ opinions $x^{k}$;
  2. (ii)
    Sensitivity mapping $\nabla_{p}h(p,d)$; and
  3. (iii)
    Gradients $\nabla_{p}\varphi(p,x)$ and $\nabla_{x}\varphi(p,x)$.
None of the above information is readily available, making a direct implementation of the recommender design (6) impractical. In fact, the users’ opinion are not accessible in practice, their dynamics are generally unknown (and so are their sensitivities), and the gradients $\nabla_{p}\varphi$, $\nabla_{x}\varphi$ depend on the users’ clicking behaviour $g(p,x)$, which is also unknown. Instead, we have access only to the users’ clicks $\{c^{k}\}$ from which we compute their CTRs $y_{i}^{k}:=\sum_{t=k-T}^{k}c_{i}^{t}/(T+1),\forall\ k\geq T$, for some time window $T$. This can be regarded as an approximation of the expectation of a click, i.e., $\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}=g(p,x)$.
To cope with these challenges, we augment the controller (6) with three auxiliary blocks associated with users’ opinions, sensitivity, and clicking behaviour estimates in real-time. Specifically, we structure the design on three levels:
  1. (i)
    Real-time opinions and users’ clicking behaviour estimation via supervised learning;
  2. (ii)
    Online sensitivity learning via Kalman filtering; and
  3. (iii)
    Gradient estimation via a forward difference method.
The resulting control architecture is illustrated in Fig. 2 and summarized in Algorithm 1. In the following three sections, we describe each level.
[…]
Figure 2: Block diagram of the proposed recommender system design with the three levels. Level I: Opinion and clicking behaviour estimation, Level II: Sensitivity estimation, Level III: Optimization.
3.1 Level I: Opinions & Clicking Behaviour Estimation
The opinions and clicking behaviours are estimated using an Artificial Neural Network (ANN). Note that the use of the ANN allows us to take into account the dynamics of opinions thus circumventing the problem of getting access to online state measurements.
We estimate the opinions by relying on implicit feedback from the users, i.e. the CTR $y\approx\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}$ of the users on recommendations with position $p$. Thus, we state that there exists a map to extract the opinions from the clicks and the positions, $x=\beta(y,p)$. To bound the resulting opinions and clicking behaviour estimation error from the ANN, we work under the following regularity assumptions [17, 39].
Assumption 3 (Opinion & Clicking behaviour maps)
We assume that
  1. i)
    the continuous map $\beta:[0,1]^{n}\times[-1,1]^{n}\to[-1,1]^{n}$ is such that $\beta(y,p)=x+\theta(x)$ with $\lVert\theta(x)\rVert\leq\theta_{x}$, for all $x\in\mathcal{X}\subseteq[-1,1]^{n}$, and $\theta_{x}<\infty$. Moreover, the image $\beta(\mathcal{Y},\mathcal{P})$ is compact for any $\mathcal{Y}\subseteq[0,1]^{n},\mathcal{P}\subseteq[-1,1]^{n}$;
  2. ii)
    there exists a constant $\alpha_{y}<\infty$ such that the composite function $g(p,\beta(y,p))=y+\nabla_{x}g(p,x)^{\top}\theta(x)+\alpha(y)$, with $\lVert\alpha(y)\rVert\leq\alpha_{y}$, for all $y\in\mathcal{Y}$. $\square$
Assumption 3 i) guarantees the map $\beta(y,p)$ to provide an estimate on users’ opinions upper bounded by $\theta_{x}$. Similarly, Assumption 3 ii), guarantees the (opinion) map $\beta(y,p)$ to (implicitly) provide an estimate on clicking behaviour upper bounded by $M_{x}\theta_{x}+\alpha_{y}$, with $M_{x}$ resulting from Assumption 2. We will qualitatively describe how these upper bound errors can be tightened in Section 3.4.
In the remainder of this section, we denote by $\mathcal{X}_{p},\mathcal{X}_{x}\in[-1,1]^{n\times m},\mathcal{X}_{y}\in[0,1]^{n\times m}$ the data sets of the positions, steady-state opinions, and the CTR used for training of the ANN. The process to acquire the training data is explained in detail in Appendix 5.1.
3.1.1 Opinion Estimation
Here, we describe the training process to estimate ${\beta}$ in Assumption 3 using an ANN.
For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p}\\ \mathcal{X}_{y,i}\end{bmatrix}\in\begin{Bmatrix}[-1,1]^{n\times m}\\ [0,1]^{1\times m}\end{Bmatrix}$, $\mathcal{X}_{x,i}\in[-1,1]^{1\times m}$, respectively.
Given that users interact with each other, the recommendations to users connected to user $i$ will indirectly influence user $i$’s opinion $x_{i}$. However, $x_{i}$ is not directly affected by the CTR of other users. To model this, we design the input layer with $n+1$ neurons for each user: $n$ neurons associated with $p$ and one neuron associated with $y_{i}$. The output layer comprises a single neuron to represent $x_{i}$. Additionally, we incorporate an intermediate layer with $n+2$ neurons. The intermediate layer employs a hyperbolic tangent activation function $\tanh(x):\mathbb{R}\rightarrow[-1,1]$, while the output layer uses an identity activation function.
Next, we provide an upper bound for the opinion estimation error $e_{x}:=h(p,d)-\hat{\beta}(y,p)$, where $\hat{\beta}$ is the opinion estimate map learned by the neural network.
Lemma 2 (Opinion estimation error)
Under Assumption 3, the opinion estimation error is upper bounded in the $\ell_{2}$-norm as $\lVert e_{x}\rVert\leq\epsilon_{x}$, with
$\epsilon_{x}:=\sqrt{n}\Big{[}$ $3\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}+2\underset{i\in[n]}{\mathrm{sup}}\omega_{\beta_{i}}(\gamma_{x})+$
$\gamma_{x}\underset{i\in[n]}{\mathrm{sup}}|v_{0,i}|\Big{]}+\theta_{x},$ (8)
where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}$ represents the maximum opinion estimation error during training of the ANN, $\omega_{\beta_{i}}(\gamma_{x})$ denotes the minimum modulus of continuity of $\beta_{i}$ on the training set as in Definition 1, and $v_{0,i}$ represents the bias from the hidden layer to the output layer of the ANN on user $i$.
Proof 3.1.
Refer to Appendix 5.2.
3.1.2 Clicking Behaviour Estimation
{tac}
For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively. We then construct an ANN similarly as for opinion estimation [^1] We omit the details due to space limitations and refer the interested reader to for a detail discussion.. {arxiv} To estimate the clicking behaviour, we proceed similarly as for opinion estimation. For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively. Given the recommender system’s personalization of positions to user’s interests, the $i$-th user’s CTR $y_{i}$ only depends on their steady-state opinion $x_{i}$ and the provided position $p_{i}$. Hence, we have two input neurons ($x_{i},p_{i}$) and one neuron in the output layer ($y_{i}$). Further, we use $3$ intermediate layers with $5$ neurons. The hyperbolic tangent and the identity are used as activation functions at the intermediate layers and the output layer, respectively.
We now derive an upper bound on the resulting clicking behaviour estimation error $e_{y}:=\hat{g}(p,\hat{x})-g(p,h(p,d))$, where $\hat{x}$ and $\hat{g}$ represents the opinion and clicking behaviour learned by the ANN.
Lemma 3.2 (Clicking behaviour estimation error).
Under Assumptions 23, the clicking behaviour estimation error is upper-bounded in the $\ell_{2}$-norm as $\lVert e_{y}\rVert\leq\epsilon_{g}$, with
$\epsilon_{g}\leq\sqrt{n}\Big{[}$ $3\underset{p\in\mathcal{X}_{p},x\in\mathcal{X}_{x}}{\mathrm{sup}}\lVert g(p,x)-\hat{g}(p,x)\rVert_{\infty}+2\underset{i\in[n]}{\mathrm{sup}}\omega_{g_{i}}(\gamma_{y})+$
$\gamma_{y}\underset{i\in[n]}{\mathrm{sup}}|w_{0,i}|\Big{]}+M_{x}\epsilon_{x}+\alpha_{y},$ (9)
where $\underset{p\in\mathcal{X}_{p},x\in\mathcal{X}_{x}}{\mathrm{sup}}\lVert g(p,x)-\hat{g}(p,x)\rVert_{\infty}$ is the maximum estimation error during ANN training, $\omega_{g_{i}}(\gamma_{y})$ denotes the minimum modulus of continuity of $g_{i}$ on the training set as in Definition 1, and $w_{0,i}$ represents the estimated bias weight from the hidden layer to the output layer of the ANN on user $i$.
{tac}
Proof 3.3.
The proof is similar to that of Lemma 2 and hence is omitted due to space limitation. We refer the interested reader to the extended version of this paper .
{arxiv}
Proof 3.4.
Refer to Appendix 5.3.
The values $\gamma_{x},\gamma_{y}$ and the modulus of continuity are properties of the training set and can be determined before training the ANN. The maximum training error term in Lemmas 2, 3.2 can be ascertained after training the ANN. The values $v_{0,i},w_{0,i}$ are the bias weights from the penultimate to the final layer in the ANN, which are known after training the ANN.
3.2 Level II: Online Sensitivity Learning
To estimate the sensitivity $\nabla_{p}h(p,d)$ in (7), we develop a Kalman filter, similar to [16]. Given that opinion dynamics and recommendations evolve over time scales that are comparable in magnitude, in order to make our OFO scheme [40] applicable, we introduce an auxiliary mechanism that guarantees time-scale separation between the opinion dynamics and the recommender system, see Remark 1.
Note that the input-output sensitivity brings information about interpersonal influence among agents, in fact, if $\nabla_{p}h_{ij}(p,d)\neq 0$ it means that the position provided to agent $j$ influences agent $i$’s opinion, hence $j$ and $i$ are connected over the platform. Thus, by estimating the sensitivity we are able to perform network-aware recommendations. We will see later in this manuscript how this turns out to be beneficial in attenuating polarization.
3.2.1 Kalman Filter for Sensitivity Learning
We denote by $\ell\in\mathbb{R}^{n^{2}}$ the vectorized sensitivity $\nabla_{p}h(p,d)\in\mathbb{R}^{n\times n}$, namely, $\ell:=\textrm{vec}(\nabla_{p}h(p,d))$, and model the sensitivity dynamics as the following random walk:
$\ell^{k}=\ell^{k-1}+w^{k-1},$ (10)
where, $w^{k}\sim\mathcal{N}(0_{n^{2}},Q^{k})$ is the process noise, with $Q^{k}$ as its corresponding covariance matrix. The measurement model is
$\Delta x_{\rm ss}^{k+1,k}=\Delta\tilde{p}^{k,k-1}\ell^{k}+v^{k},$ (11)
where $\Delta x_{\rm ss}^{k+1,k}:=h(p^{k},d)-h(p^{k-1},d)$ is the change in steady-state opinions for a corresponding change in positions $\Delta p^{k,k-1}:=p^{k}-p^{k-1}$, and $\Delta\tilde{p}^{l,m}:=\Delta(p^{l,m})^{\top}\otimes I_{n}\in\mathbb{R}^{n\times n^{2}}$, where $\otimes$ indicates the Kronecker product. The measurement noise is described by $v^{k}\sim\mathcal{N}(0_{n},R^{k})$, where $R^{k}$ is its corresponding covariance matrix. This noise accounts for the contribution of the external influence $d$ to the change of opinions $\Delta x$.
Remark 3.5.
The measurement model (11) is obtained from a first-order Taylor approximation of the steady-state map around $p^{k-1}$, i.e. $h(p^{k},d)\approx h(p^{k-1},d)+\nabla_{p}h(p^{k-1},d)^{\top}(p^{k}-p^{k-1})$. With this intuition, it can be seen that the measurement model is a linear time-variant approximation of the temporal evolution of the sensitivity $\nabla_{p}h(p^{k},d)$. It is to be noted that in reality, the steady-state $h(p^{k},d)$ cannot be perfectly achieved. Thus, it is common practice to use $\Delta x^{k+1,k}_{ss}=x^{k+1}-x^{k}$ as the measurement in (11), see [16].
The posterior update of sensitivity estimates ^$\hat{\ell}^{k}$ and covariance $\Sigma^{k}$ are given by
^$\hat{\ell}^{k}$ =^$=\hat{\ell}^{k-1}+\zeta^{k}K^{k-1}\big{(}\Delta\hat{x}^{k+1,\tau_{i}+1}-\Delta\tilde{p}^{k,\tau_{i}}\hat{\ell}^{k-1}\big{)}$
$\Sigma^{k}$ $=\Sigma^{k-1}+\zeta^{k}\big{(}Q^{k}-K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\Sigma^{k-1}\big{)},$ (12)
where $\zeta$ enforces an auxiliary trigger mechanism, with
$\zeta^{k}=\Bigg{\{}\begin{array}[]{ll}1,&k=iT\\ 0,&\rm otherwise\end{array}.$
The trigger mechanism is introduced for the recommender system to take action at stationarity, see Remark 1, and to collect a sufficient number of clicks to ensure a proper estimate of the CTR needed for opinion estimation. The trigger mechanism is turned on at integer multiples $i\in\mathbb{N}$ of the time period $T$. We define $\mathcal{T}$ as the set of all trigger time instances, with $\tau_{i}\in\mathcal{T}$ representing the most recent trigger time instance prior to $k$. Note that the update (3.2.1) uses the opinion estimate $\hat{x}$ introduced in §3.1.1 rather than the real users’ opinion, with $\hat{x}^{k+1}=\hat{\beta}(y^{k},p^{k})$, where $y^{k}:=\sum_{t=\tau_{i}}^{k}c^{t}/(k-\tau_{i}+1)$ is the CTR since the most recent trigger time instant. Finally, the Kalman filter gain $K^{k}$ in (3.2.1) is given by
$K^{k}=\Sigma^{k}(\Delta\tilde{p}^{k,\tau_{i}})^{\top}(R^{k}+\Delta\tilde{p}^{k,\tau_{i}}\Sigma^{k}(\Delta\tilde{p}^{k,\tau_{i}})^{\top})^{-1},$ (13)
with $\Delta\tilde{p}^{k,\tau_{i}}$ and $R^{k}$ as in the measurement model (11). Next, we postulate some regularity assumptions for the process and measurement models.
Assumption 4 (Gaussian noise)
The process and measurement noise $w,v$ are white Gaussian. Moreover, the steady-state opinion estimation error $e_{x}=h(p,d)-\hat{\beta}(y,p)$ is uncorrelated with $w$ and $v$. $\square$
The assumption of white noise for the process model is standard in the context of sensitivity learning in feedback optimization [16]. Intuitively, the process perturbation determines the degree of trust one puts on the sensitivity estimates. It is also reasonable to assume that the external influence is uncorrelated among users in certain cases, for example the extended FJ model in (2) where $d\equiv x_{0}$. We also note that since $\hat{x}$ is estimated using the ANN, the steady-state estimation error $e_{x}$ is not correlated with the process or measurement noises in (10), (11). Under Assumption 4, the covariances simplify as $Q^{k}=(\sigma_{q}^{k})^{2}I_{n^{2}}$ and $R^{k}=(\sigma_{r}^{k})^{2}I_{n}$, for some $\sigma_{q}$, $\sigma_{r}>0$. The tuning of $\sigma_{q},\sigma_{r}$, is based on heuristics and is described in Appendix 5.4.
To ensure that the sensitivity is correctly inferred, we must guarantee that the input positions $\Delta p$ are persistently exciting (see Definition 2). This translates to $\Delta p$ being sufficiently variable in order to explore the underlying user opinion dynamics. To do so, we work under the following assumption.
Assumption 5 (Persistency of excitation)
The inputs $\Delta p$ are persistently exciting (see Definition 2) and the trigger mechanism in (3.2.1) is activated at least $S+1$ times, namely $|\mathcal{T}|\geq S+1$. $\square$
3.3 Level III: Gradient Estimation & Optimization
We describe the gradient estimation method for the engagement cost function in (3). We remark that this algorithmic step is needed as the clicking behaviour is unknown. We then describe the OFO update rule from (6), incorporating gradient, sensitivity and state estimation.
3.3.1 Gradient Estimation
To estimate the gradient of the engagement maximization cost $\varphi^{\rm ctr}$, we use the finite forward difference method [41]
$\nabla_{x}\hat{\varphi}^{\rm ctr}_{i}(p,x)=\frac{\hat{\varphi}^{\rm ctr}(p,x+\mu e_{i})-\hat{\varphi}^{\rm ctr}(p,x)}{\mu},$ (14)
$\nabla_{p}\hat{\varphi}^{\rm ctr}_{i}(p,x)=\frac{\hat{\varphi}^{\rm ctr}(p+\mu e_{i},x)-\hat{\varphi}^{\rm ctr}(p,x)}{\mu},$ (15)
where $\nabla_{j}\hat{\varphi}^{\rm ctr}_{i},j\in\{p,x\}$ denotes the $i$-th entry of the gradient, $e_{i}\in\mathbb{R}^{n}$ refers to the $i$-th vector of the canonical basis of $\mathbb{R}^{n}$, and $\mu$ is a smoothing parameter. The cost $\hat{\varphi}^{\rm ctr}(p,x)=-1_{n}^{\top}\hat{g}(p,x)$ is evaluated as described in §3.1.2.
The smoothing parameter $\mu$ is chosen small enough so that the gradient estimate provides a good approximation of the true value. However, choosing $\mu$ in (14)–(15) too small can lead to numerical instability, as formalized in the following.
Lemma 3.6 (Gradient estimation error).
Under Assumptions 23, the gradient estimation error is upper bounded as
$\lVert\nabla_{j}\hat{\varphi}^{\rm ctr}-\nabla_{j}\varphi^{\rm ctr}\rVert\leq\frac{1}{2}L_{x}\mu+2\frac{\sqrt{n}\epsilon_{g}}{\mu},\quad j\in\{x,p\}.$
The upper bound attains its lowest value $2n^{3/4}\sqrt{\epsilon_{g}L_{j}}$ at $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}}$, with $j\in\{x,p\}$.
Proof 3.7.
Refer to Appendix 5.5.
3.3.2 Projected Gradient Descent Algorithm
We now present the projected-gradient update in (6) augmented with sensitivity, state, and gradient estimations. The augmented update reads compactly as
$p^{k+1}=\Pi_{[-1,1]^{n}}\Big{[}p^{k}-\zeta^{k}\eta\,\hat{\Phi}^{k}\Big{]},$ (16)
where the gradient surrogate $\hat{\Phi}$ is constructed by combining sensitivity, opinion, and gradient estimates as
$\hat{\Phi}^{k}=$ $\nabla_{p}\hat{\varphi}^{\rm ctr}(p^{k},\hat{x}^{k+1})+(\hat{H}^{k})^{\top}\nabla_{x}\hat{\varphi}^{\rm ctr}(p^{k},\hat{x}^{k+1})$
$+\gamma(\hat{H}^{k})^{\top}\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})-w_{\rm pe}^{k}/\eta,$ (17)
where $\hat{H}^{k}:=\nabla_{p}\hat{h}(p^{k},d)$ is the sensitivity estimate at time $k$. The additional term $w_{\rm pe}^{k}\sim\mathcal{N}(0_{n},\sigma_{\rm pe}^{2}I_{n})$ is a dither signal that ensures persistency of excitation of the inputs (Assumption 5).
In Algorithm 1, we summarize the pseudo-code of the proposed recommender system design. In the offline phase, training of the neural network is carried out for opinion and clicking behaviour estimation. In the online phase, a new sensitivity estimate is generated and new positions are provided to the users periodically, every $T$ time steps.
Algorithm 1 $\textrm{Recommender}[T,n]$
  Initialization
  Collect data using Algorithm 2
  Build opinion and clicking behaviour estimators ($\hat{\beta},\hat{g}$)
  Optimization phase
  for $k\geq 0$ do
     Collect clicks $c_{i}^{k}\sim\mathcal{B}\big{(}g_{i}(p_{i}^{k},x_{i}^{k})\big{)}$ from users
     CTR $y^{k}\leftarrow\frac{\sum_{t=\tau_{i}}^{k}c^{t}}{k-\tau_{i}+1},\tau_{i}=(i-1)T<k$
     Estimate opinions $\hat{x}_{i}^{k+1}\leftarrow\hat{\beta}_{i}(y_{i}^{k},p^{k})$
     if $\zeta^{k}=1$ then
        $\mathcal{T}\leftarrow\textrm{append}[k]$
        Estimate sensitivity $\hat{H}^{k}$ using (3.2.1)–(13)
        Estimate gradient using (14)–(15)
        Update positions $p^{k+1}$ using (16)
     else
        $\hat{H}^{k}\leftarrow\hat{H}^{k-1}$
        $p^{k+1}\leftarrow p^{k}$
     end if
  end for
3.4 Closed-loop Convergence Guarantees
We provide convergence guarantees for the sensitivity estimation process (3.2.1) and for the closed-loop interconnection between the opinions and recommendations. Before stating the closed-loop convergence theorems, we make some important observations regarding opinion estimation.
With our problem setup, the measurement information at each time instant (i.e. click/ no click) is not rich enough to produce accurate real-time opinion estimation at every time instant. As shown in Algorithm 1, opinion estimation is carried out using the CTR recorded on recommendations with a constant position since the previous trigger instant. Hence, in light of Assumption 1, the resulting opinion estimate at the end of each trigger period is an estimate of the steady-state map $h(p^{k},d)$, thus differing from real-time state estimation [17]. By considering a longer trigger period $T$, the CTR at the end of a period turns into a better approximation of $\mathbb{E}\big{[}g(p,x)\big{]}$. We qualitatively state that the modelling errors $\theta_{x},\alpha_{y}$ in Lemma 2,3.2 reduces on increasing $T$. Furthermore, as the steady-state opinions $h(p^{k},d)$ are directly estimated at each trigger instant, we circumvent the need to separately prove exponential convergence for varying inputs on the opinion dynamics.
Now, we state the main convergence result with respect to the sensitivity estimation error.
Theorem 3.8 (Sensitivity estimate convergence).
Under Assumptions 1, 4, and 5, the sensitivity estimation error $e^{k}:=\ell^{k}-\hat{\ell}^{k}$ has bias and variance bounded in norm, i.e., there exist positive constants $c_{1},c_{2},C_{f}<\infty$ and $\xi\in(0,1)$ such that
$\lVert\mathbb{E}[e^{k}]\rVert$ $\leq 2K_{m}\epsilon_{x}/(1-c_{1}\xi^{c_{2}T})+(c_{1}\xi^{c_{2}T})^{|\mathcal{T}|}\lVert\mathbb{E}[e^{0}]\rVert$
$\mathbb{E}\Big{[}\lVert e^{k}\rVert^{2}\Big{]}$ $\leq C_{f}+(c_{1}\xi^{c_{2}T})^{2|\mathcal{T}|}\mathbb{E}\Big{[}\lVert e^{0}\rVert^{2}\Big{]}$
where $C_{f}=\frac{1}{1-(c_{1}\xi^{c_{2}T})^{2}}\big{[}(T-1)\overline{\sigma}_{q}^{2}+K_{m}^{2}\big{(}\overline{\sigma}_{r}^{2}+2\epsilon_{x}^{2}\big{)}\big{]}$ with $\overline{\sigma}_{q}=\underset{k\in\mathbb{N}_{0}}{\mathrm{sup}}{\sigma}_{q}^{k}$, $\overline{\sigma}_{r}=\underset{k\in\mathbb{N}_{0}}{\mathrm{sup}}{\sigma}_{r}^{k}$, $K_{m}=\underset{k\in\mathbb{N}_{0}}{\mathrm{sup}}\lVert K^{k}\rVert$, $c_{1}\xi^{c_{2}T}<1$.
Proof 3.9.
Refer to Appendix 5.6.
Note that both bias and variance upper bounds depend on the opinion estimation error $\epsilon_{x}$. Further, note that increasing the sampling period $T$ reduces the bias and variance error through the term $1-c_{1}\xi^{c_{2}T}$. This is expected as increasing $T$ guarantees greater time-scale separation between opinion dynamics and recommendations. Finally, the upper bound on the sensitivity error variance is proportional to the noise variance of the opinions through the term $\overline{\sigma}_{r}^{2}$.
To quantify performance on the recommendations, we use the fixed-point residual mapping [42, Eq. (5)]
$\mathcal{G}(p):=\frac{1}{\eta}\Big{(}p-\Pi_{[-1,1]^{n}}\big{[}p-\eta\Phi(p,h(p,d))\big{]}\Big{)},$ (18)
The fixed-point residual mapping is zero at a critical point of (5), and is a common metric to quantify convergence of iterative algorithms in non-convex regimes [22].
The following theorem shows convergence of the closed-loop system, combining Levels 1,2,3 in Figure 2.
Theorem 3.10 (Closed-loop scheme convergence).
Let Assumptions 1-5 hold. For $\eta\in(0,\frac{1}{2L^{\prime}})$, $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}},j\in\{p,x\}$, the sequence $\{p^{k}\}_{k\in\mathbb{N}}$ generated by (16) satisfies
$\frac{1}{|\mathcal{T}|}\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}\Big{[}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$ $\leq K_{1},\quad\forall k\geq T$ (19)
where $K_{1}$ is given by
$K_{1}=\frac{6}{1-2\eta L^{\prime}}\Bigg{\{}\underset{\kappa_{1}}{\underbrace{\frac{\varphi(p^{0},h(p^{0},d))-\varphi^{*}}{3\eta|\mathcal{T}|}}}+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\Bigg{[}\underset{\kappa_{2}}{\underbrace{\gamma^{2}L^{2}\epsilon_{x}^{2}}}+$
$\underset{\kappa_{3}}{\underbrace{n^{3/2}\epsilon_{g}\!(L_{p}\!\!+\!\!L^{2}L_{x})}}\!\Bigg{]}\!\!\!+\!\!{2(2n\gamma^{2}\!\!+\!\!M_{x}^{2}\!\!+\!\!4n^{3/2}\epsilon_{g}L_{x})}\frac{\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert e^{l}\rVert^{2}]}{|\mathcal{T}|}\!\!\Bigg{\}},$
with $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$ denoting the Lipschitz constant of the gradient $\Phi(\cdot,h(\cdot,d))$, and $\varphi^{*}~{}=~{}\underset{p\in[-1,1]^{n}}{\mathrm{inf}}\varphi(p,h(p,d))~{}>~{}-n$ the cost function value at a locally optimal solution.
Proof 3.11.
Refer to Appendix 5.14.
In Theorem 3.10, we use the expectation of the second moment on the fixed-point residual mapping as the convergence measure. This metric reflects the average performance of Algorithm 1. Moreover, we note that the term $\lVert\mathcal{G}(p^{k})\rVert$ approaches zero as $p^{k}$ approaches a locally optimal solution [43]. Note that an upper bound on the minimum of $\lVert\mathcal{G}(p^{k})\rVert$ can be inferred from (19), through the inequality $\mathbb{E}\Big{[}\underset{l\in\mathcal{T}}{\mathrm{min}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}\leq\mathbb{E}\Big{[}\frac{1}{|\mathcal{T}|}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}=\frac{1}{|\mathcal{T}|}\mathbb{E}\Big{[}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$, see [19, Remark 5]. Also, note that the theoretical bound in (19) is computed only over the trigger time instances. Since the positions between two consecutive trigger time instances do not change, the term $\lVert\mathcal{G}(p^{k})\rVert$ remains the same. Thus, in order to analyze convergence, it is sufficient to compute the fixed-point residual only at the trigger time instances.
The bound $K_{1}$ shows that achieving an accurate solution of (5), i.e., a local minima $(p^{*},h(p^{*},d))$, is limited by the deviation of the initial cost at $(p^{0},h(p^{0},d))$ from the optimal one $\varphi^{*}$ (i.e. term $\kappa_{1}$), the polarization and engagement gradient estimation errors (i.e. terms $\kappa_{2},\kappa_{3}$, respectively) including both opinion and clicking behaviour estimation errors, the variance of the dither signal $\sigma_{\rm pe}^{2}$, and the sensitivity estimation error variance $\mathbb{E}[\lVert e^{l}\rVert^{2}]$.
Corollary 3.12 (Asymptotic convergence).
As $|\mathcal{T}|\rightarrow\infty$, the term $K_{1}$ approaches $K_{2}$, given by
[…]
$4(\kappa_{2}+\kappa_{3})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}\Bigg{\}}$
From $K_{2}$, we observe that the achievement of an asymptotically accurate solution of (5) at a local minima $(p^{*},h(p^{*},d))$ is limited by a combination of gradient estimation errors that already include opinion and clicking behaviour estimation (i.e. terms $\kappa_{2},\kappa_{3}$) and sensitivity estimation error, along with the variance of the dither signal $\sigma_{\rm pe}^{2}$. The term $\kappa_{1}$ approaches zero asymptotically. Moreover the upper-bound of the average of the variance of sensitivity estimation error over the trigger time instances approaches $C_{f}$ asymptotically.
4 Numerics
4.1 Opinion Dynamics and Clicking Behaviour
We assume the opinions of the users on the platform evolve according to the FJ model as in Example 1, i.e.,
$x^{k+1}=(I_{n}-\Gamma_{p}-\Gamma_{d})Ax^{k}+\Gamma_{p}p^{k}+\Gamma_{d}d.$
Further, we consider users with two different clicking behaviours: confirmation bias towards extreme positions, $C_{A}$, defined as in Example 2, and confirmation bias towards any position, $C_{B}$, wherein the click event is given by
$c_{i}\sim\mathcal{B}\Big{(}\frac{1}{2}+\frac{1}{2}e^{-4(x_{i}-p_{i})^{2}}\Big{)}.$ (20)
Note that the choice of these clicking behaviors make the steady-state optimization problem (5) non-convex.
As a proof of concept, we consider a platform with $n=15$ users, and randomly generate the adjacency matrix as $A\sim\mathcal{U}[0,1]^{n\times n}$, $A\mathbf{1}_{n}\leq\mathbf{1}_{n}$ and their initial opinions as $x^{0}\sim\mathcal{U}[-1,1]^{n}$. Moreover, we set $\Gamma_{d}=0.5\mathrm{diag}[|x^{0}|]$. This choice of $\Gamma_{d}$ indicates that the users who hold initial extreme opinions tend to be more stubborn. The external influence $d$ is modelled as a prejudice term defined as the initial opinion $x^{0}$. Further, $\Gamma_{p}$ is drawn from $\mathcal{D}_{[10^{-2},0.5]}^{n}$ so that all users are affected by the recommendations. To incorporate diversity in clicking behaviours, we randomly assign $8$ users to follow clicking behaviour $C_{A}$ and the remaining ones follow $C_{B}$.
4.2 Experimental Setting and Hyperparameter Selection
In the offline phase (Algorithm 2 in Appendix 5.1), we train the neural network for opinion and clicking behaviour estimation with $m=375$ training and $125$ testing points, simulation time $N=100$ and trigger period $T=60$, with the clicks being recorded in the duration $k\in[N-T,N]$. We set the trade-off factor $\gamma=1$ in (3), thus giving equal importance to engagement maximization and polarization mitigation.
In the online phase, we choose $\sigma_{\text{pe}}=0.07$ in (16) for the first $N/5$ time steps to facilitate sensitivity estimation and network learning, and $\sigma_{\text{pe}}=0$ afterwards to favor accurate recommendations. In the Kalman filter (3.2.1), we initialize ^$\hat{\ell}^{0}=0.5{\rm vec}(I_{n})$ and $\Sigma^{0}=I_{n^{2}}$. For gradient estimation in (14)–(15), we set the smoothing parameter $\mu=0.1$. While tuning the process noise standard deviation, we set the tuning parameter $M$ in (21) as $M=10$. Further, we start with neutral recommendations, i.e., $p^{0}=0_{n}$, and consider a periodic trigger with period $T=60$. For the polarization cost $\varphi^{\rm pol}$ in (4), we set $\epsilon_{1}=-0.5$ and $\epsilon_{2}=0.5$. All the simulations are conducted for $N=2.5\cdot 10^{4}$ time instances and over $50$ Monte-Carlo trials, with each simulation taking around 2 seconds. Since clicks are stochastic, the sequence of clicks varies across each trial. Moreover, the dither signal $w_{\rm pe}$ sequence varies across each trial, thus leading to different sensitivity estimates.
4.3 Architecture Comparisons
In this section, we compare Algorithm 1 against other similar OFO approaches that benefit from more information.
Table 2: Methods for comparison
Method Sensitivity Opinions Clicking behaviour
$M_{1}$ (Oracle)
$M_{2}$
$M_{3}$
$M_{4}$ (Alg. 1)
In Table 2, we summarize all the methods used for comparison and their attributes. In particular, for the oracle (method $M_{1}$), we employ the standard projected-gradient controller (6), wherein opinions can be directly measured, and the sensitivity and users’ clicking behavior is known.
Method $M_{2}$ uses online sensitivity estimation. Method $M_{3}$ uses opinion as well as sensitivity estimation. Finally, method $M_{4}$ corresponds to Algorithm 1, and includes also clicking behaviour and gradient estimation. All the methods’ architecture can be extracted from Figure 2 by replacing or removing some of the blocks. Method $M_{1}$ would not require the ANN in Level-1, the Kalman filter in Level-2 and gradient estimation in Level-3. Method $M_{2}$ would not require the ANN and gradient estimation. Method $M_{3}$ would need the ANN only for opinion estimation, but does not require gradient estimation.
Figure 3 shows that convergence is empirically established in all the methods through the fixed-point residual mapping norm $\lVert\mathcal{G}(p^{k})\rVert^{2}$. In particular, for methods that involve sensitivity estimation (i.e., $M_{2}-M_{4}$), we observe a higher fluctuation for the first $N/5$ time instances due to the effect of the dither signal $\sigma_{\text{pe}}$. Our proposed method offers slightly inferior performances compared to the other OFO methods. However, it must be taken into account that method $M_{4}$ is agnostic about the analytical form of $g(p,x)$, and is completely data-driven.
In Figure 4, we illustrate the relative sensitivity estimation error. It is expected that method $M_{2}$ performs better than $M_{3},M_{4}$ as the opinions are known in real time. However, the methods that involve steady-state opinion estimation, $M_{3},M_{4}$, still show comparable performance. This is due to accurate opinion estimation, as can be seen from Figure 5, where we show the opinion estimates for four randomly chosen users, along with their true values. Finally, Figure 6, shows the effect of $\gamma$ in (3) on the trade-off between engagement maximization and polarization mitigation. As expected, increasing $\gamma$, turns into a polarization cost reduction. However, the loss in engagement is not significant. This is due to the fact that roughly half of the agents follow clicking behaviour $C_{B}$, and hence still get engaged by the non-extreme content that promotes polarization mitigation. A more significant engagement loss would be expected if all the users were to be engaged by extreme content only [10].
Figure 3: Evolution of the fixed-point residuals $\lVert\mathcal{G}(p^{k})\rVert^{2}$ for the algorithms in Table 2. The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
Figure 4: Illustration of sensitivity estimation error in methods $M_{2}-M_{4}$. The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
Figure 5: Illustration of opinion estimation in method $M_{4}$.
Figure 6: Illustration of the influence of $\gamma$ in (3) on final engagement and polarization costs.
Figure 7: Steady-state mean CTR (a) and polarization (b) obtained by Algorithm 1 (OFO), its network-agnostic version (naive OFO), and the recommender design in [10]. The black lines represent the initial ideal mean CTR (left) and the initial polarization (right).
Figure 8: Evolution of the fixed-point residuals $\lVert\mathcal{G}(p^{k})\rVert^{2}$ obtained by applying Algorithm 1 and its network-agnostic version. The solid lines represent the mean and the shaded region (±$\pm 1$ standard deviation) the range of changes across the $50$ Monte Carlo trials.
4.4 Impact of Network Aware Recommendations
We showcase the impact of providing network-aware recommendations compared to individual, decoupled recommendations. To achieve this, we compare our approach with two network-agnostic methods: (i) the recommender system described in [10], and (ii) a naive version of our OFO-based recommender design.
The design in [10] provides recommendations with exclusively extreme positions (i.e., $p_{i}=\pm 1$) to each user independently. The position is calculated based on the user’s individual clicking history. Specifically, if a user more frequently clicks on $+1$ (or $-1$) positions, the system will provide $+1$ (or $-1$) positions with high probability. Conversely, if the user shows an equal likelihood of clicking on both $+1$ and $-1$ positions, the recommendations will be random.
The latter is a network-agnostic version of our recommender system design summarized in Algorithm 1. Specifically, for each user $i\in[n]$, we use a local ANN for opinion estimation that is trained using local positions and acceptance ratio only, i.e., $\hat{x}_{i}=\hat{\beta}_{i}(y_{i},p_{i})$ rather than $\hat{\beta}_{i}(y_{i},p)$. Further, we do not carry out sensitivity estimation and instead use a constant diagonal sensitivity $\hat{H}\in\mathcal{D}_{(0,0.5]}^{n}$ in (17). The rationale behind this design choice is that if the network did not contribute to engagement maximization and/or polarization mitigation, a diagonal $\hat{H}$ would yield the same performance as our network-aware approach.
We compare the performance of our recommender system with the other decoupled methods in terms of final mean CTR $J:=\sum_{k=N-T+1}^{N}\lVert c^{k}\rVert_{1}/(nT)$ and final polarization cost $\varphi^{\rm pol}(x^{N})$ in Figure 7 and the fixed point residual mapping in Figure 8. In Figure 7, we observe that our proposed algorithm achieves a slightly higher user engagement compared to its naive implementation and results in a lower polarization cost. Thus, network-aware recommendations effectively reduce polarization while maintaining high user engagement. This performance improvement is further illustrated in Figure 8, which depicts the temporal evolution of the fixed-point residuals. This result is unsurprising since polarization is a global network metric, whereas engagement maximization can be addressed individually by tailoring recommended content for each user.
The recommender design in [10], which employs only extreme recommendations, naturally leads to higher polarization costs, as shown in Figure 7(b). Furthermore, it results in lower user engagement compared to both our algorithm and the naive method. This is because some users in the network are assigned clicking behavior $C_{B}$, for which extreme recommendations do not necessarily increase engagement. Also, the setup in [10] does not consider polarization reduction as an objective.
5 Conclusion
Our aim is to improve the understanding of real-world phenomena using a simplified yet insightful model. We designed a recommender system that simultaneously maximizes user engagement and mitigates polarization. Our recommender system solely relies on clicks and does not require any prior knowledge about opinion dynamics and users’ clicking behavior. We provided theoretical optimality and closed-loop guarantees for the resulting recommender-social network interconnection. Finally, our simulations demonstrated that our recommender performs favorably against other approaches that do not leverage information about users’ interconnections. We provide evidence that polarization risk should be considered at the recommendation level.
Future research directions include relaxing the smoothness assumptions on the clicking behaviour and incorporate other interest attractors towards recommendation other than confirmation bias, e.g. repulsion. Finally, we aim for our network-centered perspective to enhance the existing literature on opinion polarization caused by algorithmic systems and inspire effective mitigation strategies.
Appendix
5.1 Training Data
The training of the neural network is carried out offline via feed-forward and back-propagation [44]. Algorithm 2, provides the pseudo-code to acquire the training data.
Algorithm 2 [$\mathcal{X}_{p},\mathcal{X}_{x},\mathcal{X}_{y}$] = $\textrm{Training}[N,T,m]$
  for $j=1$ to $m$ do
     $\overline{p}\sim\mathcal{U}[-1,1]^{n}$
     $x^{k+1}=f(x^{k},\overline{p},d^{k})$, $k\in\{0,1,\ldots,N-1\},$
     $c_{i}^{k}=\mathcal{B}(g_{i}(\overline{p}_{i},x_{i}^{k}))$, $\forall i\in[n],k\in\{0,1,\ldots,N-1\}$
     Obtain rating from users $x^{N}$
     CTR $y^{N}=\frac{1}{T+1}\sum_{k=N-T}^{N}c^{k}$
     Positions set $\mathcal{X}_{p}\leftarrow\textrm{append}[\overline{p}]$
     Steady-state opinions set $\mathcal{X}_{x}\leftarrow\textrm{append}[x^{N}]$
     CTR set $\mathcal{X}_{y}\leftarrow\textrm{append}[y^{N}]$
  end for
For each training sample $k\in[m]$, we provide news expressing a fixed random position $\overline{p}$ to the users for a period of $N$ time steps. The term $T$ is a design parameter, with $(N-T)$ representing a time instant at which opinions have reached a steady-state. After $N$ steps, we obtain the users’ rating $x^{N}$ and compute their CTR for the position $\overline{p}$ based on the last $T$ time steps. This process is carried out over $m$ Monte Carlo trials, thus collecting $m$ training data points. We emphasize that, similarly to what happens in real life with recommender systems asking the users to rate their experiences, for each trial, we only require access to one sample from the user opinion dynamics, $x^{N}$. The real time opinions in each trial $\{x^{k}\}_{k=0}^{N-1}$ are not needed.
5.2 Proof of Lemma 2
The proof follows from [45, Corollary 5.2] with the continuous function $\beta$ playing the same role as $f$ in [45].
In order to provide an explicit upper-bound for $e_{x}$ we point out that the neural network makes use of three layers: the injection layer, whose map is represented by $u:\mathbb{R}^{n+1}\to\mathbb{R}^{n+2}$, the intermediate layer whose map is represented by $z:\mathbb{R}^{n+2}\to\mathbb{R}^{n+2}$ and the output layer, whose map is represented by $v:\mathbb{R}^{n+2}\to\mathbb{R}$, with $v(z)=v^{\top}z+v_{0}$. Given that the input and output layers are linear maps, we can directly make use of [46, Theorem 7] to state the following upper-bound $\forall i\in[n]$:
$|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|\leq$ $3\underset{y_{i}\in\mathcal{X}_{y,i},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y,p)-\hat{\beta}_{i}(y,p)|+$
$2\omega_{\beta_{i}}(\gamma_{x})+|v_{0,i}|\gamma_{x},$
where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y,p)-\hat{\beta}_{i}(y,p)|$ refers to the maximum training error on each user $i$, $\omega_{\beta_{i}}(\gamma_{x})$ refers to the minimum modulus of continuity of $\beta_{i}$ on the training set, $v_{0,i}$ represents the estimated bias weight $v_{0}$ from the intermediate layer to the output layer of the neural network for user $i$. Including the modelling error $\theta_{x}$ of $\beta$ from Assumption 3, it is now possible to state the following:
$\lVert e_{x}\rVert\leq\sqrt{n}\Big{[}$ $3\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}+$
$2\underset{i\in[n]}{\mathrm{sup}}\omega_{\beta_{i}}(\gamma_{x})+\gamma_{x}\underset{i\in[n]}{\mathrm{sup}}|v_{0,i}|\Big{]}+\theta_{x}.$
{arxiv}
5.3 Proof of Lemma 3.2
Before making use of [45, Corollary 5.2] to prove the upper-bound on $e_{y}$, it is to be noted that the arguments of $g$ and $\hat{g}$, in the definition of $e_{y}$, are different. Thus, we re-write $e_{y}$ as $e_{y}=\hat{g}(p,\hat{x})-g(p,\hat{x})+g(p,\hat{x})-g(p,h(p,d))$ and note the following:
$\lVert e_{y}\rVert$ $\underset{(a)}{\leq}\lVert\hat{g}(p,\hat{x})-{g}(p,\hat{x})\rVert+\lVert g(p,h(p,d)+e_{x})-g(p,h(p,d))\rVert$
$\underset{(b)}{\leq}\lVert\hat{g}(p,\hat{x})-{g}(p,\hat{x})\rVert+\lVert\nabla_{x}g^{\top}(p,h(p,d))e_{x}\rVert+\alpha_{y}$
$\underset{(c)}{\leq}\lVert\hat{g}(p,\hat{x})-{g}(p,\hat{x})\rVert+M_{x}\epsilon_{x}+\alpha_{y}.$
In $(a)$, we make use of the identity $\hat{x}=h(p,d)+e_{x}$, with the state estimation error $e_{x}$ defined in Lemma 2. In $(b)$, we use the Taylor series expansion on $g(p,h(p,d)+e_{x})$ around $x$. In addition, the higher-order terms in the expansion of $g(p,h(p,d)+e_{x})$ are upper-bounded by the modelling error on the clicking behaviour, i.e. $\rVert\mathcal{O}(g(p,x))\lVert\leq\alpha_{y}$. In $(c)$, we use Assumption 2, i.e. $g(p,x)$ is $M_{x}$-Lipshitz with respect to $x$ and the fact that the opinion estimation error is upper-bounded with $\lVert e_{x}\rVert\leq\epsilon_{x}$ from Lemma 2.
The existence of an upper-bound on $\lVert\hat{g}(p,\hat{x})-{g}(p,\hat{x})\rVert$ is similar to the one in Lemma 2. Thus, we directly make use of [46, Theorem 7] to state the following upper-bound $\forall i\in[n]$:
$|g_{i}(p_{i},\hat{x}_{i})-\hat{g}_{i}(p_{i},\hat{x}_{i})|\leq$ $3\underset{p_{i}\in\mathcal{X}_{p,i},x_{i}\in\mathcal{X}_{x,i}}{\mathrm{sup}}|g_{i}(p_{i},x_{i})-\hat{g}_{i}(p_{i},x_{i})|+$
$2\omega_{g_{i}}(\gamma_{y})+|w_{0,i}|\gamma_{y},$
where $\underset{p_{i}\in\mathcal{X}_{p,i},x_{i}\in\mathcal{X}_{x,i}}{\mathrm{sup}}|g_{i}(p_{i},x_{i})-\hat{g}_{i}(p_{i},x_{i})|$ refers to the maximum training error on each user $i$, $\omega_{g_{i}}(\gamma_{y})$ refers to the modulus of continuity of $g_{i}$ on $\mathcal{Y}$, $w_{0,i}$ represents the estimated bias weight $w_{0}$ from the penultimate layer to the output layer of the neural network for user $i$. It is now possible to state the following:
$\lVert e_{y}\rVert\leq\sqrt{n}\Big{[}$ $3\underset{p\in\mathcal{X}_{p},x\in\mathcal{X}_{x}}{\mathrm{sup}}\lVert g(p,x)-\hat{g}(p,x)\rVert_{\infty}+$
$2\underset{i\in[n]}{\mathrm{sup}}\omega_{g_{i}}(\gamma_{y})+\gamma_{y}\underset{i\in[n]}{\mathrm{sup}}|w_{0,i}|\Big{]}+M_{x}\epsilon_{x}+\alpha_{y}$
5.4 Hyper-parameter Tuning
In order to tune the process noise covariance $Q^{k}$ we rely on heuristics and qualitative observations. Empirical results on the FJ model suggests that the sensitivity matrix is non-negative. This leads to the condition $\ell^{k}+w^{k}\geq 0_{n^{2}},\forall k\in\mathbb{N}_{0}$ in the random walk (10). Therefore, to ensure a non-negative sensitivity, it is required to have the condition $\lVert w^{k}\rVert_{\infty}\leq\mathrm{min}\{\ell^{k}\}$. This condition is satisfied by choosing an $M$ large enough such that
$\sigma_{q}^{k}=\underset{i:\hat{\ell}_{i}^{k}\neq 0}{\mathrm{min}}\hat{\ell}_{i}^{k}/M.$ (21)
Here, $M$ is a tuning parameter, to be selected ad hoc. Note in fact, that setting $M$ arbitrarily large, the bound $\lVert w^{k}\rVert_{\infty}\leq\mathrm{min}\{\ell^{k}\}$ becomes looser, but this gives more trust in the process model, resulting in poor learning performances in the posterior estimates.
The measurement noise $v$, is tuned by maximizing the measurement likelihood as in [47, Section 3D]. To this end, we solve the following optimization problem online:
$(\sigma_{r}^{k})^{2}=\underset{(\sigma_{r})^{2}}{\mathrm{arg}}{\mathrm{max}}\ P\Big{(}\Delta{x}_{\rm ss}^{k+1,\tau_{i}+1}\ |\ \Delta\tilde{p}^{k,\tau_{i}}\Big{)}$ (22)
The intuition behind (22) is to find the noise characteristics of $v$ that best represent the measurement model in (11) using the input $\Delta p$ and output $\Delta{x}$ information. Note that, since the true measurements $\Delta x$ are unavailable, they are replaced with $\Delta\hat{x}$.
5.5 Proof of Lemma 3.6
Due to Assumption 2 and Lemma 1, for any $p,x_{1},x_{2}\in\mathbb{R}^{n}$ it holds that
$\varphi^{\mathrm{ctr}}(p,x_{1})-\varphi^{\mathrm{ctr}}(p,x_{2})-(\nabla_{x}\varphi^{\mathrm{ctr}})^{\top}(p,x_{2})(x_{1}-x_{2})\leq\frac{1}{2}L_{x}\lVert x_{1}-x_{2}\rVert^{2}.$
In the above equation, we replace $x_{1},x_{2}$ with $x+\mu e_{i}$ and $x$, respectively, so that
$\varphi^{\mathrm{ctr}}(p,x+\mu e_{i})-\varphi^{\mathrm{ctr}}(p,x)-\mu(\nabla_{x}\varphi^{\mathrm{ctr}})^{\top}(p,x)e_{i}\leq\frac{1}{2}L_{x}\mu^{2}.$ (23)
We now add and subtract $\hat{\varphi}^{\mathrm{ctr}}(p,\hat{x}+\mu e_{i})$ and $\hat{\varphi}^{\mathrm{ctr}}(p,\hat{x})$ in the above equation and we note that $\hat{\varphi}^{\mathrm{ctr}}(p,\hat{x}):=-1_{n}^{\top}\hat{g}(p,\hat{x})$ and $\varphi^{\mathrm{ctr}}(p,x):=-1_{n}^{\top}g(p,x)$. Using these definitions in (23), we obtain:
$\hat{\varphi}^{\mathrm{ctr}}(p,\hat{x}+\mu e_{i})$ $-\hat{\varphi}^{\mathrm{ctr}}(p,\hat{x})\leq 1_{n}^{\top}[e_{y}(p,x)+e_{y}(p,x+\mu e_{i})]$
$+\mu(\nabla_{x}\varphi^{\mathrm{ctr}}(p,x))^{\top}e_{i}+\frac{1}{2}L_{x}\mu^{2},$ (24)
where $e_{y}(p,x)$ is the clicking behaviour estimation error as defined in (9). Dividing both sides by $\mu$ in (24) and using the gradient estimate definition in (14), we obtain
$\Big{(}\nabla_{x}\hat{\varphi}^{\mathrm{ctr}}\Big{)}_{i}(p,\hat{x})$ $-\Big{(}\nabla_{x}{\varphi}^{\mathrm{ctr}}\Big{)}_{i}(p,x)\leq\frac{L_{x}\mu}{2}+$
$\frac{1}{\mu}1_{n}^{\top}(e_{y}(p,x)+e_{y}(p,x+\mu e_{i})).$
for all $i\in[n]$. Taking the modulus on both sides, we obtain $|1_{n}^{\top}(e_{y}(p,x)+e_{y}(p,x+\mu e_{i}))|\leq\sqrt{n}\epsilon_{g}$ using the Cauchy-Schwartz inequality and the upper-bound on the clicking behaviour estimation error $\lVert e_{y}\rVert\leq\epsilon_{g}$ as in (9). We now write
$\Big{|}\Big{(}\nabla_{x}\hat{\varphi}^{\mathrm{ctr}}\Big{)}_{i}(p,\hat{x})-\Big{(}\nabla_{x}{\varphi}^{\mathrm{ctr}}\Big{)}_{i}(p,x)\Big{|}\leq\frac{1}{2}L_{x}\mu+2\frac{\sqrt{n}\epsilon_{g}}{\mu},$
from which (14) follows. Analogous reasoning is followed for the gradient estimation error with respect to $p$. To obtain the tight upper-bound, we use first-order optimality conditions with respect to $\mu$ on the term $\frac{1}{2}L_{x}\mu+2\frac{\sqrt{n}\epsilon_{g}}{\mu}$, thus obtaining $\mu^{*}=2n^{1/4}\sqrt{\epsilon_{g}/L_{x}}$. Since the second-order derivative of the term $\frac{1}{2}L_{x}\mu+2\frac{\sqrt{n}\epsilon_{g}}{\mu}$ is strictly positive, $\mu^{*}$ is the smoothing parameter that provides the lowest upper-bound on the gradient estimation error.
5.6 Proof of Theorem 3.8
The Kalman filter is uniformly asymptotically stable provided the pairs $(I_{n^{2}},\sqrt{Q^{k}}),(I_{n^{2}},\Delta\tilde{p}^{k,\tau_{i}})$ are uniformly completely controllable and observable, respectively [48, Theorem 7.4]. By design (see Appendix 5.4), the process noise covariance $Q^{k}$ is positive definite $\forall k\in\mathbb{N}_{0}$, from which the matrix pair $(I_{n^{2}},\sqrt{Q^{k}})$ is uniformly completely controllable.
To prove observability of $(I_{n^{2}},\Delta\tilde{p}^{k,\tau_{i}})$, we consider the following observability Gramian matrix [48, Chapter 7.5]:
$W_{O}(l_{1},l_{2}):=\sum_{l=l_{1}}^{l_{2}}(\Delta\tilde{p}^{\tau_{i-l+1},\tau_{i-l}})^{\top}(R^{\tau_{i-l+1}})^{-1}(\Delta\tilde{p}^{\tau_{i-l+1},\tau_{i-l}}).$ (25)
The matrix pair $(I_{n^{2}},\Delta\tilde{p}^{k,\tau_{i}})$ is said to be uniformly completely observable if $\exists\ \alpha,\beta>0$ and an $N\in\mathbb{N}_{0}:0<\alpha I_{n^{2}}\leq W_{O}(k-N,k)\leq\beta I_{n^{2}}$. Since $\Delta\tilde{p}^{t_{1},t_{2}}:=\Delta p^{t_{1},t_{2}}\otimes I_{n}$, we write $W_{O}(l_{1},l_{2})=\sum_{l=l_{1}}^{l_{2}}(\Delta{p}^{\tau_{i-l+1},\tau_{i-l}})(\Delta{p}^{\tau_{i-l+1},\tau_{i-l}})^{\top}\otimes(R^{\tau_{i-l+1}})^{-1}$. With an appropriate choice of $M$ in (21), we have $\alpha_{1}I_{n}\leq R^{k}\leq\alpha_{2}I_{n},\forall k\in\mathbb{N}_{0}$, with $\alpha_{1}>0$ and $\alpha_{2}<\infty$, leading to $\frac{1}{\alpha_{2}}\leq(R^{\tau_{i-l+1}})^{-1}\leq\frac{1}{\alpha_{1}}$. Moreover, the persistency of excitation assumption (see Assumption 5) on the inputs $\Delta p$ enable us to state that $\exists\ S\in\mathbb{N}_{0}:\beta_{1}I_{n}\leq\sum_{l=k-S}^{k}(\Delta{p}^{\tau_{i-l+1},\tau_{i-l}})(\Delta{p}^{\tau_{i-l+1},\tau_{i-l}})^{\top}\leq\beta_{2}I_{n}$, with $\beta_{1}>0$ and $\beta_{2}=(S+1)\underset{l\in[k-S,k]}{\mathrm{sup}}\ \lVert\Delta{p}^{\tau_{i-l+1},\tau_{i-l}}\rVert^{2}\leq 4n(S+1)$. This condition along with the inequality on $R^{k}$ leads to the result $0<\frac{\beta_{1}}{\alpha_{2}}I_{n^{2}}\leq W_{O}(k-S,k)\leq\frac{\beta_{2}}{\alpha_{1}}I_{n^{2}}$. Thus, the matrix pair $(I_{n^{2}},\Delta\tilde{p}^{k,\tau_{i}})$ is uniformly completely observable and it is now possible to state that the Kalman filter is uniformly asymptotically stable.
We can now derive an explicit analytical expression for the upper-bound on the bias $\lVert\mathbb{E}[e^{k}]\rVert$ and the variance $\mathbb{E}[\lVert e^{k}\rVert^{2}]$. By making use of (10) and (3.2.1) on $\ell^{k}$ and ^$\hat{\ell}^{k}$, by adding and subtracting $\Delta x_{\rm ss}^{k+1,\tau_{i}+1}$and by making use of (10) and (11) on $\Delta x_{\rm ss}^{k+1,\tau_{i}+1}$, one gets
$e^{k}$ $=\big{(}I_{n^{2}}-\zeta^{k}K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\big{)}(e^{k-1}+w^{k-1})-\zeta^{k}K^{k-1}v^{k}$
$+\ \zeta^{k}K^{k-1}\Delta e_{x}^{\tau_{i}+1,k+1}.$ (26)
Considering a trigger event at $k$, i.e. $\zeta^{k}=1$, we trace (26) back to the previous trigger instant at $\tau_{i}$, thus obtaining:
$e^{k}=$ $\big{(}I_{n^{2}}-K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\big{)}e^{\tau_{i}}+\big{(}I_{n^{2}}-K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\big{)}w^{k-1}+$
$\sum_{t=\tau_{i}}^{k-2}w^{t}-K^{k-1}v^{k}+K^{k-1}\Delta e_{x}^{\tau_{i}+1,k+1}.$ (27)
Taking the expectation on both sides of the above equation and using Assumption 4, we have:
$\mathbb{E}[e^{k}]=\big{(}I_{n^{2}}-K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\big{)}\mathbb{E}[e^{\tau_{i}}]+K^{k-1}\mathbb{E}[\Delta e_{x}^{\tau_{i}+1,k+1}].$ (28)
Since we have already established uniform complete observability, the Kalman filter gain $K^{k}$ settles to a steady-state value $K_{s}$ as $k\to\infty$. Thus, we can state that $\exists\ K_{m}:\lVert K^{k}\rVert\leq K_{m},\forall\ k\in\mathbb{N}^{0}$. Further, denoting $\psi(\tau_{i+1},\tau_{i}):=I_{n^{2}}-K^{\tau_{i+1}-1}\Delta\tilde{p}^{\tau_{i+1},\tau_{i}}$ as the state transition matrix for the Kalman filter error $e^{k}$, uniform asymptotic stability leads to $\lVert\psi(\tau_{i+1},\tau_{i})\rVert\leq c_{1}\xi^{c_{2}T}$ for some $c_{1},c_{2}>0,\xi\in(0,1):c_{1}\xi^{c_{2}T}<1$, where $\tau_{i+1}=k$ is the new trigger time instant and $T$ is the time period. Thus, taking the norm on both sides in (28), we have $\lVert\mathbb{E}[e^{\tau_{i+1}}]\rVert\leq c_{1}\xi^{c_{2}T}\lVert\mathbb{E}[e^{\tau_{i}}]\rVert+2K_{m}\epsilon_{x}$, where we note that $\mathrm{sup}\ \lVert\mathbb{E}[\Delta e_{x}^{\tau_{i}+1,k+1}]\rVert\leq\mathbb{E}[\lVert\Delta e_{x}^{\tau_{i}+1,k+1}\rVert]\leq 2\epsilon_{x}$ from Lemma 2. Hence, we derive the bias equation
$\lVert\mathbb{E}[e^{\tau_{i+1}}]\rVert\leq\big{(}c_{1}\xi^{c_{2}T}\big{)}^{|\mathcal{T}|}\lVert\mathbb{E}[e^{0}]\rVert+2K_{m}\epsilon_{x}\frac{1-(c_{1}\xi^{c_{2}T})^{|\mathcal{T}|}}{1-c_{1}\xi^{c_{2}T}},$
where, $\mathcal{T}:=\{\tau_{0},\tau_{1},\dots,\tau_{i}\}$ is the set of all trigger time instances upto $\tau_{i}$. Thus, we have the asymptotic result on the bias $\lim_{|\mathcal{T}|\to\infty}\mathbb{E}[e^{\tau_{i}}]\leq 2K_{m}\epsilon_{x}/(1-c_{1}\xi^{c_{2}T})$.
We now derive the upper-bound on the variance $\mathbb{E}[\lVert e^{k}\rVert^{2}]$. Taking the expectation of the norm squared on both sides in (26), we obtain:
$\mathbb{E}[\lVert e^{k}\rVert^{2}]\underset{(a)}{=}\mathbb{E}[\lVert$ $(I_{n^{2}}-\zeta^{k}K^{k-1}\Delta\tilde{p}^{k,\tau_{i}})(e^{k-1}+w^{k-1})-$
$\zeta^{k}K^{k-1}\big{(}v^{k}+\Delta e_{x}^{k+1,\tau_{i}+1}\big{)}\rVert^{2}]$
$\underset{(b)}{\leq}\lVert I_{n^{2}}$ $-\zeta^{k}K^{k-1}\Delta\tilde{p}^{k,\tau_{i}}\rVert^{2}(\mathbb{E}[\lVert e^{k-1}\rVert^{2}]+(\sigma_{q}^{k-1})^{2})$
+$\lVert$+ ∥ $\zeta^{k}K^{k-1}\rVert^{2}\big{(}(\sigma_{r}^{k})^{2}+2\epsilon_{x}^{2}\big{)}$
In $(b)$, we expand the norm and use Assumption 4 to state that the expectation on the cross-coupled terms are all zero. We then use $\mathbb{E}[\lVert w^{k-1}\rVert^{2}]=(\sigma_{q}^{k-1})^{2}$ and $\mathbb{E}[\lVert v^{k}\rVert^{2}]=(\sigma_{r}^{k})^{2}$. Further, Lemma 2 allows us to state that $\mathbb{E}[\lVert\Delta e_{x}^{k+1,\tau_{i}+1}\rVert^{2}]\leq 2\epsilon_{x}^{2}$. Considering a trigger at time $k$, we have:
$\mathbb{E}[\lVert e^{\tau_{i+1}}\rVert^{2}]\leq$ $\lVert I_{n^{2}}-K^{\tau_{i+1}-1}\Delta\tilde{p}^{\tau_{i+1},\tau_{i}}\rVert^{2}\mathbb{E}[\lVert e^{\tau_{i}}\rVert^{2}]+$
$(T-1)\overline{\sigma}_{q}^{2}+\lVert K^{\tau_{i+1}-1}\rVert^{2}\big{(}\overline{\sigma}_{r}^{2}+2\epsilon_{x}^{2}\big{)},$ (29)
where $\overline{\sigma}_{q}^{2}=\underset{t\in\mathbb{N}_{0}}{\mathrm{sup}}({\sigma}_{q}^{t})^{2}$ and $\overline{\sigma}_{r}^{2}=\underset{t\in\mathbb{N}_{0}}{\mathrm{sup}}({\sigma}_{r}^{t})^{2}$ with $\sigma_{q},\sigma_{r}$ estimated using (21) and (22), respectively. Tracing back (5.6) recursively to $k=0$, we obtain the following relation:
$\mathbb{E}[\lVert e^{\tau_{i+1}}\rVert^{2}]\leq$ $(c_{1}\xi^{c_{2}T})^{2|\mathcal{T}|}\mathbb{E}[\lVert e^{0}\rVert^{2}]+$
$\frac{1-(c_{1}\xi^{c_{2}T})^{2|\mathcal{T}|}}{1-c_{1}\xi^{c_{2}T}}\big{[}(T-1)\overline{\sigma}_{q}^{2}+K_{m}^{2}\big{(}\overline{\sigma}_{r}^{2}+2\epsilon_{x}^{2}\big{)}\big{]},$
We now have the following asymptotic result on the variance:
$\lim_{|\mathcal{T}|\to\infty}\mathbb{E}[\lVert e^{\tau_{i}}\rVert^{2}]\leq\frac{1}{1-c_{1}\xi^{c_{2}T}}\big{[}(T-1)\overline{\sigma}_{q}^{2}+K_{m}^{2}\big{(}\overline{\sigma}_{r}^{2}+2\epsilon_{x}^{2}\big{)}\big{]},$
5.7 Proof of Theorem 3.10
Before describing the proof, we state the following supporting lemma and remark.
Lemma 5.13 (Projections with smooth functions [43]).
Let $y=\Pi_{[-1,1]^{n}}[x-\eta u]$ with $y,x,u\in\mathbb{R}^{n}$. Then, the following inequality holds:
$\varphi(y)\leq$ $\varphi(z)+\langle y-z,\Phi(x)-u\rangle+\Big{[}\frac{L^{\prime}}{2}-\frac{1}{2\eta}\Big{]}\lVert y-x\rVert^{2}+$
$\Big{[}\frac{L^{\prime}}{2}+\frac{1}{2\eta}\Big{]}\lVert z-x\rVert^{2}-\frac{1}{2\eta}\lVert y-z\rVert^{2},\quad\forall z\in\mathbb{R}^{n}$
where $\varphi$ is the cost function to be minimized and $\Phi$ its gradient. Further, $L^{\prime}$ and $\eta$ are the smoothness factor of $\varphi$ and the step-size of the gradient descent algorithm, respectively.
Proof 5.14.
The proof is given in [43, Lemma 2].
Remark 5.15 (Composite gradient lipschitzness).
We observe that the Lipschitz and smoothness constant for $\varphi^{\mathrm{pol}}(x)$ are $2\sqrt{n}$ and $2$, respectively. Using Assumptions 1(iii), 2, the composite gradient $\Phi(p)$ is thus $L^{\prime}-$Lipschitz with respect to $p$, where $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$. Therefore, the composite cost function $\varphi(p,h(p,d))$ is $L^{\prime}-$smooth with respect to $p$.
We now state the following gradient update in the case where the gradients, sensitivity and opinions are known:
$\overline{p}^{k+1}=\Pi_{[-1,1]^{n}}[p^{k}-\eta\Phi(p^{k},h(p^{k},d))],$
where $\Phi(p,h(p,d))$ is the composite gradient. The above will serve as a benchmark to investigate the stationarity of our algorithm. We are now in the position to prove the inequalities in (19). To do so, we use Lemma 5.13 with $y=\overline{p}^{k+1}$, $x=p^{k}$ and $u=\Phi(p^{k},h(p^{k},d))$ is the composite gradient. Choosing $z=p^{k}$ and taking the expectation on both sides, we obtain:
$\mathbb{E}\Big{[}\varphi(\overline{p}^{k+1})\Big{]}\leq\mathbb{E}\Big{[}\varphi(p^{k})+\Big{(}\frac{L^{\prime}}{2}-\frac{1}{\eta}\Big{)}\lVert\overline{p}^{k+1}-p^{k}\rVert^{2}\Big{]}.$ (30)
We now define the update step with our algorithm:
${p}^{k+1}=\Pi_{[-1,1]^{n}}[p^{k}-\eta\zeta^{k}\hat{\Phi}^{k}],$
where, the composite gradient estimate $\hat{\Phi}^{k}$ is given in (17). We use Lemma 5.13 with $y=p^{k+1}$, $x=p^{k}$ and $u=\zeta^{k}\hat{\Phi}^{k}$. Choosing $z=\overline{p}^{k+1}$ and taking the expectation on both sides of the inequality, we obtain:
$\mathbb{E}\Big{[}\varphi(p^{k+1})\Big{]}\leq\mathbb{E}\Big{[}$ $\varphi(\overline{p}^{k+1})+\Big{(}\frac{L^{\prime}}{2}-\frac{1}{2\eta}\Big{)}\lVert p^{k+1}-p^{k}\rVert^{2}+$
$\langle p^{k+1}-\overline{p}^{k+1},\Phi(p^{k},h(p^{k},d))-\zeta^{k}\hat{\Phi}^{k}\rangle+$ (31)
$\Big{(}\frac{L^{\prime}}{2}+\frac{1}{2\eta}\Big{)}\lVert\overline{p}^{k+1}-p^{k}\rVert^{2}-\frac{1}{2\eta}\lVert p^{k+1}-\overline{p}^{k+1}\rVert^{2}\Big{]}$
We now add inequalities (30) and (31) to obtain:
$\mathbb{E}\Big{[}\varphi(p^{k+1})\Big{]}\leq\mathbb{E}\Big{[}$ $\varphi(p^{k})+\Big{(}\frac{L^{\prime}}{2}-\frac{1}{2\eta}\Big{)}\lVert p^{k+1}-p^{k}\rVert^{2}+$
$\underset{T_{1}}{\underbrace{\langle p^{k+1}-\overline{p}^{k+1},\Phi(p^{k},h(p^{k},d))-\zeta^{k}\hat{\Phi}^{k}\rangle}}+$ (32)
$\Big{(}{L^{\prime}}-\frac{1}{2\eta}\Big{)}\lVert\overline{p}^{k+1}-p^{k}\rVert^{2}-\frac{1}{2\eta}\lVert p^{k+1}-\overline{p}^{k+1}\rVert^{2}\Big{]}.$
We now focus on the term $T_{1}$. Using Cauchy-Schwartz relation and the fact that the geometric mean of two non-negative real numbers is always less than its arithmetic mean, we obtain the following: $T_{1}\leq\lVert p^{k+1}-\overline{p}^{k+1}\rVert\ \lVert\Phi(p^{k},h(p^{k},d))-\zeta^{k}\hat{\Phi}^{k}\rVert\leq\frac{1}{2\eta}\lVert p^{k+1}-\overline{p}^{k+1}\rVert^{2}+\frac{\eta}{2}\lVert\Phi(p^{k},h(p^{k},d))-\zeta^{k}\hat{\Phi}^{k}\rVert^{2}.$ Using the above inequality in (32), we obtain:
$\mathbb{E}\Big{[}\varphi(p^{k+1})\Big{]}\leq\mathbb{E}\Big{[}\varphi(p^{k})+\underset{T_{2}}{\underbrace{\Big{(}\frac{L^{\prime}}{2}-\frac{1}{2\eta}\Big{)}\lVert p^{k+1}-p^{k}\rVert^{2}}}+$ (33)
$\frac{\eta}{2}{{\lVert\Phi(p^{k},h(p^{k},d))-\zeta^{k}\hat{\Phi}^{k}\rVert^{2}}}+\Big{(}{L^{\prime}}-\frac{1}{2\eta}\Big{)}\eta^{2}\lVert\mathcal{G}(p^{k})\rVert^{2}\Big{]}.$
In the above inequality, we used the definition of fixed-point residual mapping according to (18).
We now assume that there is a trigger at time instant $k$, i.e. $\zeta^{k}=1$. To obtain a feasible upper-bound on $\mathbb{E}[\lVert\mathcal{G}(p^{k})\rVert^{2}]$, we need $L^{\prime}-1/2\eta<0$. Thus, the step-size is constrained with $\eta\in(0,\frac{1}{2L^{\prime}})$. With this constraint, we have $T_{2}\leq 0$. This leads to the following:
$\mathbb{E}\Big{[}\lVert\mathcal{G}(p^{k})\rVert^{2}\Big{]}$ $\leq\frac{2}{\eta(1-2\eta L^{\prime})}\Big{\{}\mathbb{E}\Big{[}\varphi(p^{k})\Big{]}-\mathbb{E}\Big{[}\varphi(p^{k+1})\Big{]}+$
$\frac{\eta}{2}\underset{T_{3}}{\underbrace{\mathbb{E}\Big{[}\lVert\Phi(p^{k},h(p^{k},d))-\hat{\Phi}^{k}\rVert^{2}\Big{]}}}\Big{\}}.$ (34)
We now analyze the term $T_{3}$. For the sake of convenience, we drop the arguments of the gradient $p,x$ and time argument $k$ in the gradient terms. Thus, we denote $\nabla_{p}\varphi^{\rm ctr}=\nabla_{p}\varphi^{\rm ctr}(p^{k},h(p^{k},d))$, $\nabla_{x}\varphi^{\rm ctr}=\nabla_{x}\varphi^{\rm ctr}(p^{k},h(p^{k},d))$, $\nabla_{p}\hat{\varphi}^{\rm ctr}=\nabla_{p}\hat{\varphi}^{\rm ctr}(p^{k},\hat{x}^{k+1})$ and $\nabla_{x}\hat{\varphi}^{\rm ctr}=\nabla_{x}\hat{\varphi}^{\rm ctr}(p^{k},\hat{x}^{k+1})$. Using the definitions of $\Phi(p^{k},h(p^{k},d)),\hat{\Phi}^{k}$, we have:
$T_{3}=\mathbb{E}\Big{[}$ $\lVert\nabla_{p}\varphi^{\rm ctr}-\nabla_{p}\hat{\varphi}^{\rm ctr}+H^{k\top}\nabla_{x}\varphi^{\rm ctr}-\hat{H}^{k\top}\nabla_{x}\hat{\varphi}^{\rm ctr}+$ (35)
$\gamma H^{k\top}\nabla_{x}\varphi^{\rm pol}(h(p^{k},d))-\gamma\hat{H}^{k\top}\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})+\frac{w_{\rm pe}^{k}}{\eta}\rVert^{2}\Big{]},$
where $H^{k}=\nabla_{p}h(p^{k},d)$ is the true sensitivity and $\hat{H}^{k}$ is its estimate at time $k$. We now analyze the upper-bound on $T_{3}$:
[…]
$\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+\gamma^{2}\lVert(H^{k})^{\top}(\nabla_{x}\varphi^{\rm pol}(h(p^{k},d))-\nabla_{x}{\varphi}^{\rm pol}(\hat{x}^{k+1}))\rVert^{2}+$
$\big{(}\gamma^{2}\lVert\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})\rVert^{2}+\lVert\nabla_{x}\hat{\varphi}^{\rm ctr}\rVert^{2}\big{)}\mathbb{E}\Big{[}\lVert H^{k}-\hat{H}^{k}\rVert^{2}\Big{]}\Big{\}}$
[…]
$\gamma^{2}\lVert(H^{k})^{\top}(\nabla_{x}\varphi^{\rm pol}(h(p^{k},d))-\nabla_{x}{\varphi}^{\rm pol}(\hat{x}^{k+1}))\rVert^{2}+$
$\big{(}\gamma^{2}\lVert\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})\rVert^{2}+\lVert\nabla_{x}\hat{\varphi}^{\rm ctr}\rVert^{2}\big{)}\mathbb{E}\Big{[}\lVert H^{k}-\hat{H}^{k}\rVert^{2}\Big{]}\Big{\}}$
[…]
$\big{(}\gamma^{2}\lVert\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})\rVert^{2}+\lVert\nabla_{x}\hat{\varphi}^{\rm ctr}\rVert^{2}\big{)}\mathbb{E}\Big{[}\lVert H^{k}-\hat{H}^{k}\rVert^{2}\Big{]}\Big{\}}$
[…]
$\big{(}\gamma^{2}\lVert\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})\rVert^{2}+\lVert\nabla_{x}\hat{\varphi}^{\rm ctr}\rVert^{2}\big{)}\mathbb{E}\Big{[}\lVert H^{k}-\hat{H}^{k}\rVert^{2}\Big{]}\Big{\}}$
[…]
$\big{(}4n\gamma^{2}+\lVert\nabla_{x}\hat{\varphi}^{\rm ctr}\pm\nabla_{x}{\varphi}^{\rm ctr}\rVert^{2}\big{)}\mathbb{E}\Big{[}\lVert e^{k}\rVert^{2}\Big{]}\Big{\}}$
[…]
$\big{(}4n\gamma^{2}+2(M_{x}^{2}+4n^{3/2}\epsilon_{g}L_{x})\big{)}\mathbb{E}\Big{[}\lVert e^{k}\rVert^{2}\Big{]}\Big{\}}$ (36)
To obtain $(a)$, we added and subtracted $(H^{k})^{\top}\nabla_{x}\hat{\varphi}^{\rm ctr}$ and $\gamma(H^{k})^{\top}\nabla_{x}\varphi^{\rm pol}(\hat{x}^{k+1})$ inside the norm in (35). We then used the fact that $\mathbb{E}\Big{[}\lVert\sum_{j=1}^{m}r_{j}\rVert^{2}\Big{]}\leq m\sum_{j=1}^{m}\mathbb{E}\Big{[}\lVert r_{j}\rVert^{2}\Big{]}$. We also make use of the fact that $\mathbb{E}[\lVert w_{\rm pe}^{k}\rVert^{2}]=\sigma_{\rm pe}^{2}$. In $(b)$, we made use of Lemma 3.6 for the gradient estimate accuracy on $\varphi^{\rm ctr}$ and Assumption 1(iii) for Lipschitz condition on $h(p,d)$ to arrive at the term $4n^{3/2}\epsilon_{g}(L_{p}+L^{2}L_{x})$. In $(c)$, we made use of Remark 5.15 for the smoothness condition on $\varphi^{\rm pol}$ and Assumption 1(iii) for Lipschitz condition on $h(p,d)$ to arrive at the term $4L^{2}\lVert x-\hat{x}\rVert^{2}$. In $(d)$, we made use of the fact that the norm of the steady-state opinion estimation error is upper-bounded by $\epsilon_{x}$. In $(e)$, we made use of Remark 5.15 for the Lipschitz condition on $\varphi^{\rm pol}$, thus arriving at the term $4n$. We also add and subtract the term $\nabla_{x}{\varphi}^{\rm ctr}$. Further, we use inequality $\lVert H^{k}-\hat{H}^{k}\rVert\leq\lVert H^{k}-\hat{H}^{k}\rVert_{F}=\lVert e^{k}\rVert$, where $\lVert\cdot\rVert_{F}$ refers to the Frobenius norm and $e^{k}$ is the sensitivity estimation error. In $(f)$, we made use of the fact that $\lVert a+b\rVert^{2}\leq 2(\lVert a\rVert^{2}+\lVert b\rVert^{2})$. We then use Assumption 2 to state that $\lVert\nabla_{x}\varphi^{\rm ctr}\rVert\leq M_{x}$ and Lemma 3.6 to state the upper-bound on $\lVert\nabla_{x}\varphi^{\rm ctr}-\nabla_{x}\hat{\varphi}^{\rm ctr}\rVert$.
We now use the inequality (36) in (34) and add the inequalities over the trigger time instances up to $k$, leading to telescopic cancellation. Thus, we have:
$\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert\mathcal{G}(p^{l})\rVert^{2}$ $]\leq\frac{2}{\eta(1-2\eta L^{\prime})}\Big{\{}\mathbb{E}\Big{[}\varphi(p^{0})\Big{]}-\mathbb{E}\Big{[}\varphi(p^{k+1})\Big{]}\Big{\}}+$
$\frac{6|\mathcal{T}|}{1-2\eta L^{\prime}}\Big{\{}4n^{3/2}\epsilon_{g}(L_{p}+L^{2}L_{x})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\gamma^{2}L^{2}\epsilon_{x}^{2}\Big{\}}+$
$\frac{12\big{[}2n\gamma^{2}+(M_{x}^{2}+4n^{3/2}\epsilon_{g}L_{x})\big{]}}{1-2\eta L^{\prime}}\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}\Big{[}\lVert e^{l}\rVert^{2}\Big{]}.$
Since the positions do not change between two consecutive trigger time instances, it is sufficient to investigate convergence guarantees at the trigger time instances.
Using Theorem 3.8 for the upper-bound on $\mathbb{E}[\lVert e^{k}\rVert^{2}]$, the summation of this term over the trigger instances leads to the following:
$\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}\Big{[}\lVert e^{l}\rVert^{2}\Big{]}\leq$ $|\mathcal{T}|C_{f}+\mathbb{E}\Big{[}\lVert e^{0}\rVert^{2}\Big{]}\sum_{\underset{l\leq k}{l\in\mathcal{T}}}(c_{1}\xi^{c_{2}T})^{2|\mathcal{T}|}$ (37)
$|\mathcal{T}|C_{f}+\mathbb{E}\Big{[}\lVert e^{0}\rVert^{2}\Big{]}\Big{(}\frac{1-(c_{1}\xi^{c_{2}T})^{2|\mathcal{T}|}}{1-(c_{1}\xi^{c_{2}T})^{2}}\Big{)}.$
We start the algorithm with $p^{0}=0_{n}$, thus $\mathbb{E}\Big{[}\varphi(p^{0})\Big{]}=\varphi(0)$. Further, $\exists\ \varphi^{*}\leq\mathbb{E}\Big{[}\varphi(p^{k})\Big{]},\forall k\in\mathbb{N}_{0}$, i.e. $\varphi^{*}$ is a local optimal value. Thus, with the above formulations and (37), we obtain (19).

### v2

@@
-Mitigating Polarization in Recommender Systems via Network-aware Feedback Optimization
+Network-aware Recommender System via Online Feedback Optimization
 Abstract
-We consider a recommender system that takes into account the interaction between recommendations and the evolution of user interests.
-Users opinions are influenced by both social interactions and recommended content.
-We leverage online feedback optimization to design a recommender system that trades-off between maximizing engagement and minimizing polarization.
-The recommender system is agnostic about users’ opinion, clicking behavior, and social interactions, and solely relies on clicks.
-We establish optimality and closed-loop stability of the resulting feedback interconnection between the social platform and the recommender system.
-We numerically validate our algorithm when the user population follows an extended Friedkin–Johnsen model.
-We observe that network-aware recommendations significantly reduce polarization without compromising user engagement.
+Personalized content on social platforms can exacerbate negative phenomena such as polarization, partly due to the feedback interactions between recommendations and the users.
+In this paper, we present a control-theoretic recommender system that explicitly accounts for this feedback loop to mitigate polarization.
+Our approach extends online feedback optimization – a control paradigm for steady-state optimization of dynamical systems – to develop a recommender system that trades off users engagement and polarization reduction, while relying solely on online click data.
+We establish theoretical guarantees for optimality and stability of the proposed design and validate its effectiveness via numerical experiments with a user population governed by Friedkin–Johnsen dynamics.
+Our results show these “network-aware” recommendations can significantly reduce polarization while maintaining high levels of user engagement.
 {IEEEkeywords}
@@
 1 Introduction
-Problem Description
 Online social platforms use recommender systems to provide users with tailored content to maximize engagement over the platform.
-To do so, recommender systems exploit methods, such as content-based filtering [1] and collaborative filtering [2], that combine information personalization, popularity and similarity of interests with other users.
-An important drawback of content personalization is related to the amplification phenomenon [3] associated with the feedback loop that arises between recommended content and user preferences.
-Amplification leads to misinformation spread [4], filter bubbles [5], bias exacerbation [6] and ultimately polarization [7, 8].
-Driven by these considerations we propose a recommender system that takes into account the interplay between recommendations and user interests.
-Our goal is to understand i) the impact of recommendations on user opinions and ii) how should recommendation system depart from engagement maximization to mitigate polarization.
-By leveraging on online feedback optimization (OFO) [9] and building upon [10, 11], we design a network-aware, multiple objective, recommender system as a feedback controller, illustrated in Figure 1.
+To do so, recommender systems exploit techniques, such as content-based filtering [1] and collaborative filtering [2], that combine information personalization, popularity and similarity of interests with other users.
+A significant drawback of content personalization is related to the amplification phenomenon [3], associated with the feedback loop that arises between recommended content and user preferences.
+Amplification leads to misinformation spread [4], filter bubbles [5], bias exacerbation [6] and, ultimately, polarization [7, 8].
+Motivated by these challenges, we consider instead recommender systems inspired by control-theoretic principles that have the potential of explicitly accounting for the feedback interplay between recommendations and user interests to mitigate polarization.
+Our goals are to develop a systems-theory-based recommender design to understand i) the impact of recommendations on user opinions, and ii) how recommender systems should depart from engagement maximization to mitigate polarization.
+We approach the problem by building upon the systems-theory perspective of recommender systems pioneered in [9, 10] and by leveraging on online feedback optimization (OFO) [11] – an emerging control paradigm for steady-state optimization of dynamical systems.
+Specifically, we aim at designing a recommender system as a dynamic feedback controller that mitigates polarization by providing users personalized content, using solely implicit feedback from their online click data, as illustrated in Figure 1.
+Figure 1: Closed-loop interconnection between users’ opinion in a social network and the proposed recommender system.
+Related Work
+Recommender systems.
+Most of the literature on recommender systems relies on static methods (such as matrix factorization) and neglects the dynamic nature of the feedback interactions between the users and the online platforms [12].
+In contrast, our work develops along a more recent systems-theory perspective on recommender systems [9, 10], where the feedback interaction between social platform and recommender systems is made explicit.
+In [9], the recommender system provides only extreme positions to the users based on their clicking history.
+The authors show that when confirmation bias [13] – the tendency to engage with content that aligns with prior beliefs – is the sole factor driving clicks, their recommender system promotes opinion polarization.
+The recommender design proposed in [14] alternates between exploitation and exploration by either providing news which more closely align with the users’ beliefs or exploring new content.
+The authors show that shifts in the opinions at the individual level do not necessarily reflect in the opinion distribution of the population.
+In [15], the recommender system is modeled as an artificial user within the social platform, providing news content at minimum distance from all users opinions to promote engagement.
+This study investigates how such a recommender system can affect opinion dynamics across the network.
+In [10], the focus is on the interplay between a single user, the recommender system, and harmful content.
+The authors seek recommendation policies that balance maximizing the click-through rate (CTR) and mitigate harm.
+Their findings suggest that to properly mitigate harm, it is necessary to account for user dynamics.
+Notably, most of these studies focus solely on individual-user designs, where recommendations are tailored to each user in isolation, without explicitly accounting for the broader influence of social interactions.
+However, social interactions are well-known to be a crucial factor in shaping opinions over a connected population [16, 17, 18].
+Online feedback optimization.
+Most OFO controllers are based on first-order optimization algorithms [11], and their deployment require real-time access to the states of the plant, an explicit knowledge of its input-output sensitivity (or a closed approximation), and a closed-form expression of the gradient of the control objective.
+However, some recent extensions [19, 20, 21, 22, 23] have relaxed some of these stringent requirements.
+Specifically, [19] proposes to estimate the input-output sensitivity of the plant from real-time state measurements using recursive-least square; while [20] achieves similar results leveraging historical input-output trajectories and behavioral systems theory.
+In [21], the authors address the case where exact states measurements cannot be directly accessed, but need to be inferred from perceptual information.
+Their solution relies on the integration of a neural network component that transform high-dimensional sensory data into state estimates.
+Similarly, [22] proposes a neural-network solution for the case where a closed-form expression of the gradient is not available.
+Finally, [23] utilizes a zeroth-order method to estimate the composite gradient directly from online cost function evaluations, eliminating the need for both input-output sensitivity and a closed-form gradient expression.
+See Table 1 for a comparative overview of these OFO controllers.
+Table 1: Comparative overview of the main requirements of some recent online feedback optimization (OFO) controllers.
+Controller Sensitivity States Objective Gradient Convex
+[11]
+[19, 20]
+[21]
+[22, Alg. 2]
+[23]
+This work
+Contributions
+Unfortunately, existing OFO designs cannot be directly applied to social networks, as recommender systems typically only have access to implicit feedback in the form of click data.
+Real-time users opinion are inaccessible, social interactions may be hidden due to privacy preferences, and there is no closed-form expression for the expected CTR as it depends on unknown users clicking behavior – namely, their tendencies or pattern when interacting with recommended items.
+In this paper, we address these challenges by integrating auxiliary estimation blocks into a projected-gradient OFO controller to compensate for the missing information, thereby enabling online feedback optimization in social networks.
+Specifically, we estimate opinions and clicking behaviour using supervised learning, sensitivity using Kalman filtering and gradients using zero-order methods.
+Unlike most system-theory-based recommender system designs in the literature, our proposed algorithm is network-aware, namely, it leverages on users’ social interactions to better mitigate undesired population phenomena such as polarization, while maintaining high levels of engagement.
+Our contribution is threefold:
+  1. (i)
+    We propose a novel OFO-based recommender system algorithm that steers a social network towards steady-state opinion profiles that balance between maximizing overall click-through rate and minimizing polarization.
+Our design is agnostic about real-time opinions of the users, their network interactions and clicking behavior, and solely relies on click data measured online.
+  2. (ii)
+    We establish closed-loop stability guarantees of the recommender system-social network interconnection with respect to locally-optimal solutions of the steady-state control objective, assuming exponential stability of the opinion dynamics and other mild technical conditions.
+  3. (iii)
+    We validate the proposed recommender system through numerical simulations on a social network where user opinions are governed by an extended version of the Friedkin–Johnsen (FJ) dynamics [9], which incorporates the influence of recommended content.
+Our results demonstrate that network-awareness plays a crucial role in reducing opinion polarization in social platforms.
+Outline
+In Section 2, we formulate the recommender objectives as a steady-state control problem.
+In Section 3, we describe the recommender design using OFO.
+Section 4 present numerical simulations to evaluate the performance of our algorithm and highlight the benefits of network-awareness.
+Finally, Section 5 summarizes our findings and concludes the paper.
+Notations and Preliminaries
 […]
-Figure 1: Illustration of the closed-loop between the opinion dynamics and the recommender system.
-Related Work
-Recommendation systems.
-Most of the works on recommender system are static (e.g.
-matrix factorization) and neglect the dynamic nature of the feedback interaction between the user and the online platforms.
-Our work develops along a recent systems-theory perspective on recommender systems [10] where the feedback interaction between social platform and recommender systems is made explicit.
-In [10], the recommender system provides only extreme positions to the user based on its clicking history.
-They show that when confirmation bias [12], namely the tendency to engage with content that aligns with prior beliefs, is the only driving force to click on contents, the recommender system induces opinion polarization.
-The recommender design proposed in [13] alternates between exploitation and exploration by either providing news which more closely align with the users’ beliefs or exploring new content.
-They show that shifts in the opinions at the individual level do not necessarily reflect in the opinion distribution of the population.
-The work [14] models the recommender system as an artificial user in the network that provides news content at minimum distance from all users opinions to promote engagement.
-The authors study how recommender systems of this form affect opinion dynamics.
-In [11], they consider the interplay between one single user, the recommender and harmful content and seek recommendation policies that tradeoff between maximizing the click-through rate (CTR) and mitigate harm.
-Their finding suggest that to properly mitigate harm, it is necessary to account for user dynamics.
-Importantly, most of the works related to recommender systems largely focus on one single user [11, 13, 10] without explicitly considering the influence of social interactions.
-Online feedback optimization.
-Conventional OFO algorithms employ first-order gradient methods [15] that require knowledge about real-time states, sensitivity of the input-output steady-state map and the gradients of the objective function with respect to the input and the states.
-Recent advancements [16, 17, 18, 19], however allow the employment of OFO with less information.
-In particular, in [16] the sensitivity matrix is estimated online using online state information.
-In [17, 18] an OFO scheme is proposed where supervised learning is employed for real-time states and gradient estimation, respectively.
-Finally, in [19], zeroth-order methods are used to directly estimate the composite gradient using cost function evaluations, thus not requiring sensitivity estimation.
-See Table 1 for an overview of the recent advancements in OFO.
-Table 1: State-of-the-art OFO algorithms.
-OFO Method Sensitivity States Cost function Gradients Convexity
-[15]
-[16]
-[17]
-[18, Alg. 2]
-[19]
-Our work
-Contributions
-Recommender systems provide users tailored content based on implicit feedback, such as clicks.
-We design an OFO-based recommender system algorithm that extracts information on state, sensitivity and objective function gradient from clicks and dynamically provides user-tailored recommendations.
-The recommender system encompasses polarization mitigation and engagement maximization in the objective function.
-We show that by incorporating information about real-time opinions and clicking behaviors of users as well as the topology of the social network, the recommender is able to enhance engagement while mitigating polarization.
-The recommendation algorithm is interpreted as a controller that incorporates opinion (state), topology (sensitivity), and clicking behavior (gradient) estimations.
-The system dynamics is assumed to be uniformly exponentially stable for constant inputs and to admit a continuously differentiable steady-state map.
-We provide theoretical closed-loop guarantees between the opinion dynamics and the recommender.
-Numerical results show that our network-aware recommender system compares favourably in terms of polarization with negligible engagement losses against isolated user scenarios, where the recommendations are solely provided based on personalization and network interactions are neglected.
-Our contribution is threefold:
-  1. 1.
-    We propose a recommender system algorithm that takes into account users’ preferences and trades off between maximizing CTR and minimizing polarization.
-The recommender system is agnostic about real-time opinions, network interactions and clicking behaviour.
-  2. 2.
-    From a technical perspective, the OFO algorithm runs combining estimations of system state, input-output sensitivity, and objective function gradient on a non-convex optimization problem.
-As an independent technical contribution, we provide convergence guarantees towards a locally optimal solution, thereby improving the state of the art on feedback optimization.
-  3. 3.
-    We validate the proposed control scheme on synthetic data generated via an extension of the Friedkin–Johsnen (FJ) model [10] that also takes into account the recommended content influence over opinions.
-We show the importance of network awareness on the attenuation of opinion polarization.
-Outline
-In Section 2, we formulate the optimization problem for engagement maximization and polarization mitigation.
-In Section 3, we describe the recommender design using OFO.
-In Section 4, we provide numerical simulations to show our algorithm performances and illustrate the benefits of network-awareness.
-Section 5 concludes our work.
-Notations and Preliminaries
-We let the symbols $\mathbb{R}$($\mathbb{R}_{+}$), $\mathbb{N}_{0}$ denote the set of (positive) real numbers and non-negative integers, respectively.
-The set of integers $\{1,2,\ldots,n\}$ is denoted by $[n]$.
-For a vector $y\in\mathbb{R}^{n}$, we let the symbol $|y|$ denote the vector whose $i$-th entry, $|y|_{i}$, is the $i$-th component of $y$ in modulus, $|y_{i}|$.
-The symbol $\langle\cdot,\cdot\rangle:\mathbb{R}^{n}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$ denotes the standard inner product on $\mathbb{R}^{n}$.
-The value of a signal $x$ at time instant $k$ is denoted by $x^{k}$.
-The symbol $1_{n},(0_{n})$, denotes the all ones (zeros) vector of size $n$.
-The symbols $I_{n}$ and $O_{n}$, denote the $n$-dimensional identity and zero matrix, respectively.
-Given a matrix $A\in\mathbb{R}^{m\times n}$, the matrix is positive (non-negative) if all its elements are greater (or equal) than zero and we denote it as $A>0$ ($A\geq 0$).
-We let $\mathrm{vec}(A)$ denote the vectorized version of $A$, i.e., if $A=[a_{1}\ a_{2}\ \ldots\ a_{n}]$, with $a_{j}\in\mathbb{R}^{m},j\in[n]$, then $\mathrm{vec}(A)=[a_{1}^{\top}a_{2}^{\top}\ldots a_{n}^{\top}]^{\top}$.
-For a matrix $A\in\mathbb{R}^{n\times n}$, we denote by $\mathrm{det}[A]$ its determinant; $A$ is row-stochastic if $A\geq 0$ and $A1_{n}=1_{n}$.
-The symbol $\mathcal{D}_{\mathcal{U}}\subseteq\mathbb{R}$ denotes the set of $n$-dimensional diagonal matrices with diagonal entries in $\mathcal{U}$.
-The symbol $\mathrm{diag}[x]$ denotes the diagonal matrix with $x_{i}$ on its $i$-th diagonal entry.
-Given $\mathcal{C}\subset\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, we let $\Pi_{\mathcal{C}}[x]$ denote the Euclidean projection of $x$ over $\mathcal{C}$.
-The normal distribution is denoted by $\mathcal{N}(\mu,\Sigma)$, with $\mu,\Sigma$ representing the mean and variance, and the uniform distribution over the interval $[a,b]$, with $a,b,\in\mathbb{R}$ is represented by $\mathcal{U}[a,b]$.
 We now state some preliminary lemmas and definitions that are recurrently used in this manuscript.
@@
 Definition 2 (Persistently exciting input [25])
-The signal $u\in\mathbb{R}^{n}$ is said to be persistently exciting of order $n$ if there exists a time span $S>0$ such that for all $k>0$, the matrix formed by the columns $u^{k+i}$ for $i\in\{0,1,\ldots S\}$ has full rank, i.e.
-$\mathrm{Rank}[u^{k},u^{k+1},\ldots,u^{k+S}]=n$.
+The signal $u\in\mathbb{R}^{n}$ is said to be persistently exciting of order $n$ if there exists a time span $S>0$ such that for all $k>0$, the matrix formed by the columns $u^{k+i}\in\mathbb{R}^{n}$ for $i\in\{0,1,\ldots S\}$ has full rank, i.e.
+$\mathrm{Rank}[u^{k}\ u^{k+1}\ \ldots\ u^{k+S}]=n$.
 Definition 3 (Global $\beta$-smoothness [26])
-A continuously differentiable map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is globally $\beta$-smooth if its gradient $\nabla\Phi$ is $\beta$-Lipschitz.
+A continuously differentiable map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is globally $\beta$-smooth if its gradient $\nabla\Phi$ is globally $\beta$-Lipschitz.
 Lemma 1 (Properties of $\beta$-smooth functions [26])
 Given a globally $\beta$-smooth map $\Phi:\mathbb{R}^{n}\rightarrow\mathbb{R}$, then $\Phi(x_{1})-\Phi(x_{2})-\nabla\Phi^{\top}(x_{2})(x_{1}-x_{2})\leq\frac{1}{2}\beta\lVert x_{1}-x_{2}\rVert^{2}$.
-Moreover, if $\Phi$ is twice continuously-differentiable, then $\lVert\nabla^{2}\Phi\rVert\leq\beta$.
+Further, if $\Phi$ is twice continuously-differentiable, then $\lVert\nabla^{2}\Phi\rVert\leq\beta$.
 2 Problem setup
@@
     The map $h(p,d)$ is continuously-differentiable and $L-$Lipschitz [26] with respect to $p$. $\square$
-This assumption can be easily adopted for a wide array of renowned opinion dynamics models, such as De-Groot model [23], FJ model [24], and polar opinion dynamics model [25], among others, for which the influence of a recommendation system can be, in principle, included.
+This assumption can be easily adopted for a wide array of renowned opinion dynamics models, such as De-Groot model [17], FJ model [16], and polar opinion dynamics model [18], among others, for which the influence of a recommender system can be, in principle, included.
 For an extensive list of such models, we refer the reader to the survey [27].
@@
 Engagement of the users over the platform is measured in terms of CTR, measured as number of clicks over number of provided news articles.
-As common practice in the context of implicit feedback recommendation systems [32], for a generic user $i\in[n]$ holding an opinion $x_{i}$, we model the probability of $i$ clicking on a recommendation with position $p_{i}$ as a random variable $c_{i}$ sampled from a Bernoulli distribution with unknown argument $g_{i}(p_{i},x_{i})$, i.e.
+As common practice in the context of implicit feedback recommender systems [33], for a generic user $i\in[n]$ holding an opinion $x_{i}$, we model the probability of $i$ clicking on a recommendation with position $p_{i}$ as a random variable $c_{i}$ sampled from a Bernoulli distribution with unknown argument $g_{i}(p_{i},x_{i})$, i.e.
 $c_{i}\sim\mathcal{B}(g_{i}(p_{i},x_{i}))$.
@@
 Assumption 2 (Clicking behaviour)
-The clicking behaviour $g(p,x):[-1,1]^{n}\times[-1,1]^{n}\to[0,1]^{n}$ is $M_{x}$-Lipschitz with respect to $x$, and $L_{p}$- and $L_{x}$-smooth with respect to $p$ and $x$, respectively. $\square$
-A concrete example of a clicking behaviour is provided.
+The clicking behaviour $g(p,x):[-1,1]^{n}\times[-1,1]^{n}\to[0,1]^{n}$ is $M_{x}$ Lipschitz with respect to $x$, and $L_{p}$- and $L_{x}$-smooth with respect to $p$ and $x$, respectively. $\square$
+Next, we provide an example of a clicking behaviour.
 Example 2 (Extremity confirmation bias [9])
-A user $i\in[n]$, holding opinion $x_{i}$, affected by extremity confirmation bias clicks on a recommendation $p_{i}$ with probability
+A user $i~{}\in~{}[n]$, holding opinion $x_{i}$, affected by extremity confirmation bias clicks on a recommendation $p_{i}$ with probability
 $c_{i}\sim\mathcal{B}\left(\frac{1}{2}+\frac{1}{2}x_{i}p_{i}\right),$
 where the clicking behaviour $g_{i}(p_{i},x_{i})=\frac{1}{2}+\frac{1}{2}x_{i}p_{i}$ models confirmation bias towards extreme recommendations.
-In fact, if the opinion $x_{i}\approx\pm 1$ and the position $p_{i}\approx\pm 1$ ($\mp 1$), then the probability of clicking is almost $1$ (00).
+In fact, for an opinion $x_{i}\approx\pm 1$ and a position $p_{i}\approx\pm 1$ ($\mp 1$), the probability of clicking is almost $1$ (00).
 The click event becomes random as the user position approaches the neutral stance $p_{i}\approx 0$, highlighting diminished engagement for less polarized content or when recommendations directly counter the user’s stance on the issue.
@@
 The goal of the recommender system is to provide recommendations that optimize a specific metric, denoted as $\varphi(p,x)$.
-Typical choices include maximizing CTR [10, 11], opinion control [33], and polarization mitigation [34].
-Deviating from traditional recommendation systems, we consider a multi-objective cost function combining engagement maximization and polarization mitigation, given by
+Typical choices in the literature include maximizing CTR [9, 10], opinion control [34], and polarization mitigation [35].
+Deviating from traditional recommender systems, we consider a multi-objective cost function combining engagement maximization and polarization mitigation, given by
 $\varphi(p,x)=\varphi^{\text{ctr}}(p,x)+\gamma\,\varphi^{\text{pol}}(x),$ (3)
-where $\gamma$ is a non-negative trade-off parameter. The engagement-related term in the cost (3) is the CTR, formalized as
+where $\gamma$ is a non-negative trade-off parameter. The engagement-related term in (3) is the CTR, formalized as
 $\varphi^{\text{ctr}}(p,x)=-\sum_{i\in[n]}\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,],$
 where $\mathbb{E}_{c_{i}\sim\mathbb{B}(g_{i}(x_{i},p_{i}))}[\,c_{i}\,]$ is the expectation of user $i$’s clicking, given their opinion $x_{i}$ and a recommendation with position $p_{i}$.
-The second term in (3), $\varphi^{\rm pol}$, is a smooth Lipschitz function that measures opinion polarization.
-For our algorithm we consider it to be $\varphi^{\text{pol}}(x):=\sum_{i\in[n]}s_{i}(x)$, where $s_{i}$ is a soft penalty function:
+The second term in (3), $\varphi^{\rm pol}$, is a smooth and Lipschitz function that measures opinion polarization.
+For our algorithm we consider it to be $\varphi^{\text{pol}}(x):=\sum_{i\in[n]}s_{i}(x)$, where $s_{i}$ penalizes deviations from the moderate (zero) opinions:
 $s_{i}(x_{i})=\begin{cases}(x_{i}-\epsilon_{1})^{2}&x_{i}<\epsilon_{1}\\ 0&\epsilon_{1}\leq x_{i}\leq\epsilon_{2}\\ (\epsilon_{2}-x_{i})^{2}&x_{i}>\epsilon_{2}\end{cases}$ (4)
@@
 According to [36], the metric in (4) falls in the class of polarization spread, indicating how extreme the opinions are.
-Note that, for $\epsilon_{1}=\epsilon_{2}=0$, the term (4) reduces to $\varphi^{\text{pol}}(x)=\lVert x\rVert^{2}$, a classic measure for opinion polarization [28].
+Note that, for $\epsilon_{1}=\epsilon_{2}=0$, the term (4) reduces to $\varphi^{\text{pol}}(x)=\lVert x\rVert^{2}$, another classic measure for opinion polarization [29].
 The recommender system aims at finding the recommendation that drives the user population with dynamics (1) to a steady-state minimizing the multi-objective cost (3):
@@
 If the opinion dynamics in (1) and the clicking behavior $g_{i}$ were known, and an accurate prediction of external influences $d$ were available, then one could aim to solve (5) offline.
-However, in practice, none of these information is readily available, and the recommender system provides positions solely based on clicks $\{c_{i}^{k}\}$.
+However, in practice, none of these information is readily available, and the recommender system can provide personalized positions solely relying on clicks $\{c_{i}^{k}\}$.
 Remark 1
-Equation (5b) translates into the users’ opinion stationarity before the recommender system action is taken.
-This is consistent with most of the works on recommender systems based on preference dynamics[11, 10, 36, 6, 37], where the action is taken based on the steady-state dynamics of the proposed model.
+Equation (5b) reflects the assumption that users’ opinions reach steady-state before the recommender system takes action.
+This approach is in line with most of the existing works on recommender systems that incorporates preference dynamics [10, 9, 37, 6, 38], where recommendation decisions are made based on the steady-state behavior of the users’ preferences under the model.
 Our feedback control problem is formally stated as follows:
@@
 Design a feedback controller $p(c)$ so that (1) converges to a solution $(p^{*},x^{*})$ of the optimization problem (5), assuming that only clicks $c_{i}^{k}$ are available.
-The clicking behaviour $g_{i}(p_{i},x_{i})$, steady-state map $h(p,d)$, the opinion dynamics (1), external influence $d$ are unknown, and the real-time opinions $x$ are not measurable.
+The clicking behaviours $g_{i}(p_{i},x_{i})$, steady-state map $h(p,d)$, the opinion dynamics (1), external influence $d$ are unknown, and the real-time opinions $x$ are not measurable.
 3 Recommender system design
-We approach Problem 1 by designing a dynamic feedback controller inspired by the projected-gradient descent algorithm in [38, Section 3A].
+We approach Problem 1 by designing a dynamic feedback controller inspired by the projected-gradient descent algorithm in [39, Section 3A].
 The resulting recommender system dynamically generates recommendations as
@@
   1. (i)
-    Real-time opinions and users’ clicking behaviour estimation via supervised learning;
+    Estimation of real-time opinions and users’ clicking behaviour via supervised learning;
   2. (ii)
@@
 The opinions and clicking behaviours are estimated using an Artificial Neural Network (ANN).
-Note that the use of the ANN allows us to take into account the dynamics of opinions thus circumventing the problem of getting access to online state measurements.
-We estimate the opinions by relying on implicit feedback from the users, i.e.
-the CTR $y\approx\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}$ of the users on recommendations with position $p$.
-Thus, we state that there exists a map to extract the opinions from the clicks and the positions, $x=\beta(y,p)$.
+We estimate the opinions by relying on implicit feedback from the users, i.e., the CTR $y\approx\mathbb{E}\big{[}\mathcal{B}(g(p,x))\big{]}$ of the users on recommendations with position $p$.
+Thus, we seek a map to extract the opinions from the clicks and the positions, $x=\beta(y,p)$.
 To bound the resulting opinions and clicking behaviour estimation error from the ANN, we work under the following regularity assumptions [21, 40].
@@
   1. i)
-    the continuous map $\beta:[0,1]^{n}\times[-1,1]^{n}\to[-1,1]^{n}$ is such that $\beta(y,p)=x+\theta(x)$ with $\lVert\theta(x)\rVert\leq\theta_{x}$, for all $x\in\mathcal{X}\subseteq[-1,1]^{n}$, and $\theta_{x}<\infty$.
-Moreover, the image $\beta(\mathcal{Y},\mathcal{P})$ is compact for any $\mathcal{Y}\subseteq[0,1]^{n},\mathcal{P}\subseteq[-1,1]^{n}$;
+    There exists a continuous map $\beta:[0,1]^{n}\times[-1,1]^{n}\to[-1,1]^{n}$ is such that $\beta(y,p)=x+\theta(x)$ with $\lVert\theta(x)\rVert\leq\theta_{x}$, for all $x\in\mathcal{X}\subseteq[-1,1]^{n}$, and $\theta_{x}<\infty$; moreover, the image $\beta(\mathcal{Y},\mathcal{P})$ is compact for any $\mathcal{Y}\subseteq[0,1]^{n},\mathcal{P}\subseteq[-1,1]^{n}$.
   2. ii)
-    there exists a constant $\alpha_{y}<\infty$ such that the composite function $g(p,\beta(y,p))=y+\nabla_{x}g(p,x)^{\top}\theta(x)+\alpha(y)$, with $\lVert\alpha(y)\rVert\leq\alpha_{y}$, for all $y\in\mathcal{Y}$.
+    There exists a constant $\alpha_{y}<\infty$ such that the composite function takes the form $g(p,\beta(y,p))=y+\nabla_{x}g(p,x)^{\top}\theta(x)+\alpha(y)$, with $\lVert\alpha(y)\rVert\leq\alpha_{y}$, for all $y\in\mathcal{Y}$.
 $\square$
 Assumption 3 i) guarantees the map $\beta(y,p)$ to provide an estimate on users’ opinions upper bounded by $\theta_{x}$.
-Similarly, Assumption 3 ii), guarantees the (opinion) map $\beta(y,p)$ to (implicitly) provide an estimate on clicking behaviour upper bounded by $M_{x}\theta_{x}+\alpha_{y}$, with $M_{x}$ resulting from Assumption 2.
+Similarly, Assumption 3 ii), guarantees the (opinion) map $\beta(y,p)$ to (implicitly) provide an estimate on clicking behavior upper bounded by $M_{x}\theta_{x}+\alpha_{y}$, with $M_{x}$ resulting from Assumption 2.
 We will qualitatively describe how these upper bound errors can be tightened in Section 3.4.
@@
 Here, we describe the training process to estimate ${\beta}$ in Assumption 3 using an ANN.
-For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p}\\ \mathcal{X}_{y,i}\end{bmatrix}\in\begin{Bmatrix}[-1,1]^{n\times m}\\ [0,1]^{1\times m}\end{Bmatrix}$, $\mathcal{X}_{x,i}\in[-1,1]^{1\times m}$, respectively.
+For each user $i\in[n]$, the input and output training sets are
+$\begin{bmatrix}\mathcal{X}_{p}\\ \mathcal{X}_{y,i}\end{bmatrix}\in\begin{Bmatrix}[-1,1]^{n\times m}\\ [0,1]^{1\times m}\end{Bmatrix},\mathcal{X}_{x,i}\in[-1,1]^{1\times m},$
+respectively.
 Given that users interact with each other, the recommendations to users connected to user $i$ will indirectly influence user $i$’s opinion $x_{i}$.
@@
 Lemma 2 (Opinion estimation error)
-Under Assumption 3, the opinion estimation error is upper bounded in the $\ell_{2}$-norm as $\lVert e_{x}\rVert\leq\epsilon_{x}$, with
+Under Assumption 3, the opinion estimation error is upper bounded as $\lVert e_{x}\rVert\leq\epsilon_{x}$ with
 $\epsilon_{x}:=\sqrt{n}\Big{[}$ $3\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}+2\underset{i\in[n]}{\mathrm{sup}}\omega_{\beta_{i}}(\gamma_{x})+$
 $\gamma_{x}\underset{i\in[n]}{\mathrm{sup}}|v_{0,i}|\Big{]}+\theta_{x},$ (8)
-where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}$ represents the maximum opinion estimation error during training of the ANN, $\omega_{\beta_{i}}(\gamma_{x})$ denotes the minimum modulus of continuity of $\beta_{i}$ on the training set as in Definition 1, and $v_{0,i}$ represents the bias from the hidden layer to the output layer of the ANN on user $i$.
+where $\underset{y\in\mathcal{X}_{y},p\in\mathcal{X}_{p}}{\mathrm{sup}}\lVert\beta(y,p)-\hat{\beta}(y,p)\rVert_{\infty}$ represents the maximum opinion estimation error during training of the ANN, $\omega_{\beta_{i}}(\gamma_{x})$ is the minimum modulus of continuity of $\beta_{i}$ on the training set as in Definition 1, and $v_{0,i}$ denotes the bias from the hidden layer to the output layer of the ANN on user $i$.
 Proof 3.1.
@@
 {arxiv} To estimate the clicking behaviour, we proceed similarly as for opinion estimation.
-For each user $i\in[n]$, the input and output training sets are $\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m}$, $\mathcal{X}_{y,i}\in[0,1]^{1\times m}$, respectively.
-Given the recommender system’s personalization of positions to user’s interests, the $i$-th user’s CTR $y_{i}$ only depends on their steady-state opinion $x_{i}$ and the provided position $p_{i}$.
+For each user $i\in[n]$, the input and output training sets are
+$\begin{bmatrix}\mathcal{X}_{p,i}\\ \mathcal{X}_{x,i}\end{bmatrix}\in[-1,1]^{2\times m},\mathcal{X}_{y,i}\in[0,1]^{1\times m},$
+respectively.
+Given the recommender system’s personalization of positions to the user’s interests, the $i$-th user’s CTR $y_{i}$ only depends on their steady-state opinion $x_{i}$ and the provided position $p_{i}$.
 Hence, we have two input neurons ($x_{i},p_{i}$) and one neuron in the output layer ($y_{i}$).
@@
 Lemma 3.2 (Clicking behaviour estimation error).
-Under Assumptions 23, the clicking behaviour estimation error is upper-bounded in the $\ell_{2}$-norm as $\lVert e_{y}\rVert\leq\epsilon_{g}$, with
+Under Assumptions 23, the clicking behaviour estimation error is upper-bounded as $\lVert e_{y}\rVert\leq\epsilon_{g}$, with
 $\epsilon_{g}\leq\sqrt{n}\Big{[}$ $3\underset{p\in\mathcal{X}_{p},x\in\mathcal{X}_{x}}{\mathrm{sup}}\lVert g(p,x)-\hat{g}(p,x)\rVert_{\infty}+2\underset{i\in[n]}{\mathrm{sup}}\omega_{g_{i}}(\gamma_{y})+$
@@
 To estimate the sensitivity $\nabla_{p}h(p,d)$ in (7), we develop a Kalman filter, similar to [19].
-Given that opinion dynamics and recommendations evolve over time scales that are comparable in magnitude, in order to make our OFO scheme [40] applicable, we introduce an auxiliary mechanism that guarantees time-scale separation between the opinion dynamics and the recommender system, see Remark 1.
+Given that opinion dynamics and recommendations evolve over time scales that are comparable in magnitude, in order to make our OFO scheme [41] applicable, we introduce an auxiliary mechanism that guarantees time-scale separation between the opinion dynamics and the recommender system updates, see Remark 1.
 Note that the input-output sensitivity brings information about interpersonal influence among agents, in fact, if $\nabla_{p}h_{ij}(p,d)\neq 0$ it means that the position provided to agent $j$ influences agent $i$’s opinion, hence $j$ and $i$ are connected over the platform.
@@
 Remark 3.5.
-The measurement model (11) is obtained from a first-order Taylor approximation of the steady-state map around $p^{k-1}$, i.e.
-$h(p^{k},d)\approx h(p^{k-1},d)+\nabla_{p}h(p^{k-1},d)^{\top}(p^{k}-p^{k-1})$.
+The measurement model (11) is obtained from a first-order Taylor approximation of the steady-state map around $p^{k-1}$, i.e., $h(p^{k},d)\approx h(p^{k-1},d)+\nabla_{p}h(p^{k-1},d)^{\top}(p^{k}-p^{k-1})$.
 With this intuition, it can be seen that the measurement model is a linear time-variant approximation of the temporal evolution of the sensitivity $\nabla_{p}h(p^{k},d)$.
-It is to be noted that in reality, the steady-state $h(p^{k},d)$ cannot be perfectly achieved.
-Thus, it is common practice to use $\Delta x^{k+1,k}_{ss}=x^{k+1}-x^{k}$ as the measurement in (11), see [16].
+Note that, the steady-state $h(p^{k},d)$ is reached only asymptotically.
+Thus, it is common practice [19] to use instead $\Delta x^{k+1,k}_{ss}=x^{k+1}-x^{k}$ as the measurement in (11).
+$\square$
 The posterior update of sensitivity estimates ^$\hat{\ell}^{k}$ and covariance $\Sigma^{k}$ are given by
@@
 $\zeta^{k}=\Bigg{\{}\begin{array}[]{ll}1,&k=iT\\ 0,&\rm otherwise\end{array}.$
-The trigger mechanism is introduced for the recommender system to take action at stationarity, see Remark 1, and to collect a sufficient number of clicks to ensure a proper estimate of the CTR needed for opinion estimation.
+The trigger mechanism is introduced for the recommender system to update positions close to steady state, see Remark 1, and to collect a sufficient number of clicks to ensure a proper estimate of the CTR needed for opinion estimation.
 The trigger mechanism is turned on at integer multiples $i\in\mathbb{N}$ of the time period $T$.
@@
 Intuitively, the process perturbation determines the degree of trust one puts on the sensitivity estimates.
-It is also reasonable to assume that the external influence is uncorrelated among users in certain cases, for example the extended FJ model in (2) where $d\equiv x_{0}$.
+It is also reasonable to assume that the external influence is uncorrelated among users in certain cases, for example the extended FJ model in (2) where $d=x_{0}$.
 We also note that since $\hat{x}$ is estimated using the ANN, the steady-state estimation error $e_{x}$ is not correlated with the process or measurement noises in (10), (11).
@@
 The inputs $\Delta p$ are persistently exciting (see Definition 2) and the trigger mechanism in (3.2.1) is activated at least $S+1$ times, namely $|\mathcal{T}|\geq S+1$. $\square$
+In practice, this assumption can be satisfied by introducing a dither signal in the update rule (6).
 3.3 Level III: Gradient Estimation & Optimization
@@
 where $\nabla_{j}\hat{\varphi}^{\rm ctr}_{i},j\in\{p,x\}$ denotes the $i$-th entry of the gradient, $e_{i}\in\mathbb{R}^{n}$ refers to the $i$-th vector of the canonical basis of $\mathbb{R}^{n}$, and $\mu$ is a smoothing parameter.
-The cost $\hat{\varphi}^{\rm ctr}(p,x)=-1_{n}^{\top}\hat{g}(p,x)$ is evaluated as described in §3.1.2.
+Note that in the two-point gradient estimation, we do not actuate the opinion dynamics twice but rather use the engagement cost function estimate $\hat{\varphi}^{\rm ctr}(p,x)=-1_{n}^{\top}\hat{g}(p,x)$ as described in §3.1.2.
 The smoothing parameter $\mu$ is chosen small enough so that the gradient estimate provides a good approximation of the true value.
@@
 3.4 Closed-loop Convergence Guarantees
-We provide convergence guarantees for the sensitivity estimation process (3.2.1) and for the closed-loop interconnection between the opinions and recommendations.
-Before stating the closed-loop convergence theorems, we make some important observations regarding opinion estimation.
-With our problem setup, the measurement information at each time instant (i.e.
-click/ no click) is not rich enough to produce accurate real-time opinion estimation at every time instant.
-As shown in Algorithm 1, opinion estimation is carried out using the CTR recorded on recommendations with a constant position since the previous trigger instant.
-Hence, in light of Assumption 1, the resulting opinion estimate at the end of each trigger period is an estimate of the steady-state map $h(p^{k},d)$, thus differing from real-time state estimation [17].
-By considering a longer trigger period $T$, the CTR at the end of a period turns into a better approximation of $\mathbb{E}\big{[}g(p,x)\big{]}$.
-We qualitatively state that the modelling errors $\theta_{x},\alpha_{y}$ in Lemma 2,3.2 reduces on increasing $T$.
-Furthermore, as the steady-state opinions $h(p^{k},d)$ are directly estimated at each trigger instant, we circumvent the need to separately prove exponential convergence for varying inputs on the opinion dynamics.
+Here, we establish convergence guarantees for the sensitivity estimation process (3.2.1) and for the closed-loop interconnection between the opinions and recommendations.
+Before stating the formal results, we make some important observations regarding our opinion estimation technique and its implication on the closed-loop convergence analysis.
+In our problem setup, the measurement information at each time instant (i.e., a single click/no click data), is not rich enough to produce accurate estimates for the users’ real-time opinion.
+As shown in Algorithm 1, the opinion estimation is instead carried out using the CTR recorded on multiple recommendations with a constant position since the previous trigger instant.
+In light of the exponential stability of the opinion dynamics (i.e., Assumption 1), the resulting estimate at the end of each trigger period is, in practice, close to an estimate of the steady-state opinion $h(p^{k},d)$.
+This approach is therefore substantially different from real-time state estimation [21].
+Thus, as we directly estimate steady-state opinions at each trigger instant, we can greatly simplify our closed-loop stability analysis by treating the opinion dynamics as a static mapping.
+Also note that as the trigger period $T$ increases, the resulting CTR offers a better approximation of the CTR in expectation $\mathbb{E}\big{[}g(p,x)\big{]}$.
+The modeling errors $\theta_{x},\alpha_{y}$ in Lemma 23.2 exist due to a finite time horizon $T$ and reduces as $T$ increases.
 Now, we state the main convergence result with respect to the sensitivity estimation error.
@@
 Finally, the upper bound on the sensitivity error variance is proportional to the noise variance of the opinions through the term $\overline{\sigma}_{r}^{2}$.
-To quantify performance on the recommendations, we use the fixed-point residual mapping [42, Eq. (5)]
-$\mathcal{G}(p):=\frac{1}{\eta}\Big{(}p-\Pi_{[-1,1]^{n}}\big{[}p-\eta\Phi(p,h(p,d))\big{]}\Big{)},$ (18)
-The fixed-point residual mapping is zero at a critical point of (5), and is a common metric to quantify convergence of iterative algorithms in non-convex regimes [22].
-The following theorem shows convergence of the closed-loop system, combining Levels 1,2,3 in Figure 2.
-Theorem 3.10 (Closed-loop scheme convergence).
-Let Assumptions 1-5 hold. For $\eta\in(0,\frac{1}{2L^{\prime}})$, $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}},j\in\{p,x\}$, the sequence $\{p^{k}\}_{k\in\mathbb{N}}$ generated by (16) satisfies
+To quantify performance on the recommendations, we use the fixed-point residual mapping [43, Eq. (5)]
+$\mathcal{G}(p):=\frac{1}{\eta}\Big{(}p-\Pi_{[-1,1]^{n}}\big{[}p-\eta\Phi(p,h(p,d))\big{]}\Big{)}.$ (18)
+The fixed-point residual mapping is zero at any critical point of (5), and is a common metric to quantify convergence of iterative algorithms in non-convex regimes [26].
+The following theorem shows the convergence of the closed-loop system, combining Levels I, II, and III in Figure 2, using the second moment of the fixed-point residual mapping.
+Since the problem is non-convex, the trajectories of $\{p^{k}\}_{k=0}^{N}$ vary across each trial.
+Thus, it is essential to evaluate $\lVert\mathcal{G}(p^{k})\rVert$ with its expectation.
+Moreover, we compute the average of this metric over all the trigger time instances to reflect the average performance of Algorithm 1 over the time in which recommendations are provided.
+The theoretical bound is only computed over the trigger time instances since the positions remain constant between two trigger instances, and so does the residual $\lVert\mathcal{G}(p^{k})\rVert$.
+Theorem 3.10 (Closed-loop convergence).
+Let Assumptions 15 hold.
+For any $\eta\in(0,\frac{1}{2L^{\prime}})$ and $\mu^{*}=2n^{1/4}\sqrt{{\epsilon_{g}}/{L_{j}}},j\in\{p,x\}$, the sequence $\{p^{k}\}_{k\in\mathbb{N}}$ generated by (16) satisfies
 $\frac{1}{|\mathcal{T}|}\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}\Big{[}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$ $\leq K_{1},\quad\forall k\geq T$ (19)
-where $K_{1}$ is given by
-$K_{1}=\frac{6}{1-2\eta L^{\prime}}\Bigg{\{}\underset{\kappa_{1}}{\underbrace{\frac{\varphi(p^{0},h(p^{0},d))-\varphi^{*}}{3\eta|\mathcal{T}|}}}+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\Bigg{[}\underset{\kappa_{2}}{\underbrace{\gamma^{2}L^{2}\epsilon_{x}^{2}}}+$
-$\underset{\kappa_{3}}{\underbrace{n^{3/2}\epsilon_{g}\!(L_{p}\!\!+\!\!L^{2}L_{x})}}\!\Bigg{]}\!\!\!+\!\!{2(2n\gamma^{2}\!\!+\!\!M_{x}^{2}\!\!+\!\!4n^{3/2}\epsilon_{g}L_{x})}\frac{\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert e^{l}\rVert^{2}]}{|\mathcal{T}|}\!\!\Bigg{\}},$
-with $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$ denoting the Lipschitz constant of the gradient $\Phi(\cdot,h(\cdot,d))$, and $\varphi^{*}~{}=~{}\underset{p\in[-1,1]^{n}}{\mathrm{inf}}\varphi(p,h(p,d))~{}>~{}-n$ the cost function value at a locally optimal solution.
+where the upper bound $K_{1}$ is given by
+$K_{1}=\frac{6}{1-2\eta L^{\prime}}\Bigg{(}\underset{\kappa_{1}}{\underbrace{\frac{\varphi(p^{0},h(p^{0},d))-\varphi^{*}}{3\eta|\mathcal{T}|}}}+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}+4\Bigg{(}\underset{\kappa_{2}}{\underbrace{\gamma^{2}L^{2}\epsilon_{x}^{2}}}+$
+$\underset{\kappa_{3}}{\underbrace{n^{3/2}\epsilon_{g}\!(L_{p}\!\!+\!\!L^{2}L_{x})}}\!\Bigg{)}\!\!\!+\!\!{2(2n\gamma^{2}\!\!+\!\!M_{x}^{2}\!\!+\!\!4n^{3/2}\epsilon_{g}L_{x})}\frac{\sum_{\underset{l\leq k}{l\in\mathcal{T}}}\mathbb{E}[\lVert e^{l}\rVert^{2}]}{|\mathcal{T}|}\ \!\!\Bigg{)},$
+In there, $L^{\prime}=L_{p}+L^{2}(L_{x}+2)$ is the Lipschitz constant of the gradient $\Phi(\cdot,h(\cdot,d))$, and $\varphi^{*}={\mathrm{inf}}_{p\in[-1,1]^{n}}\varphi(p,h(p,d))>-n$ is the cost function value at a locally optimal solution.
 Proof 3.11.
 Refer to Appendix 5.15.
-In Theorem 3.10, we use the expectation of the second moment on the fixed-point residual mapping as the convergence measure.
-This metric reflects the average performance of Algorithm 1.
-Moreover, we note that the term $\lVert\mathcal{G}(p^{k})\rVert$ approaches zero as $p^{k}$ approaches a locally optimal solution [43].
-Note that an upper bound on the minimum of $\lVert\mathcal{G}(p^{k})\rVert$ can be inferred from (19), through the inequality $\mathbb{E}\Big{[}\underset{l\in\mathcal{T}}{\mathrm{min}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}\leq\mathbb{E}\Big{[}\frac{1}{|\mathcal{T}|}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}=\frac{1}{|\mathcal{T}|}\mathbb{E}\Big{[}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$, see [19, Remark 5].
-Also, note that the theoretical bound in (19) is computed only over the trigger time instances.
-Since the positions between two consecutive trigger time instances do not change, the term $\lVert\mathcal{G}(p^{k})\rVert$ remains the same.
-Thus, in order to analyze convergence, it is sufficient to compute the fixed-point residual only at the trigger time instances.
-The bound $K_{1}$ shows that achieving an accurate solution of (5), i.e., a local minima $(p^{*},h(p^{*},d))$, is limited by the deviation of the initial cost at $(p^{0},h(p^{0},d))$ from the optimal one $\varphi^{*}$ (i.e.
-term $\kappa_{1}$), the polarization and engagement gradient estimation errors (i.e.
-terms $\kappa_{2},\kappa_{3}$, respectively) including both opinion and clicking behaviour estimation errors, the variance of the dither signal $\sigma_{\rm pe}^{2}$, and the sensitivity estimation error variance $\mathbb{E}[\lVert e^{l}\rVert^{2}]$.
+Note that an upper bound on the minimum of $\lVert\mathcal{G}(p^{k})\rVert$ can be inferred from (19), through the inequality $\mathbb{E}\Big{[}\underset{l\in\mathcal{T}}{\mathrm{min}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}\leq\mathbb{E}\Big{[}\frac{1}{|\mathcal{T}|}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}=\frac{1}{|\mathcal{T}|}\mathbb{E}\Big{[}\sum_{l\in\mathcal{T}}\lVert\mathcal{G}(p^{l})\rVert^{2}\Big{]}$, see [23, Remark 5].
+Thus, intuitively, the bound in (19) describes the maximum possible minimum gap between the trajectory $\{p^{k},h(p^{k},d)\}_{k=0}^{N}$ obtained using Algorithm 1 and a local minima $(p^{*},h(p^{*},d))$.
+The bound $K_{1}$ shows that achieving an accurate solution of (5), i.e., a local minima $(p^{*},h(p^{*},d))$, is limited by the deviation of the initial cost at $(p^{0},h(p^{0},d))$ from the optimal one $\varphi^{*}$ (i.e., the term $\kappa_{1}$), the polarization and engagement gradient estimation errors (i.e., the terms $\kappa_{2},\kappa_{3}$, respectively) including both opinion and clicking behaviour estimation errors, the variance of the dither signal $\sigma_{\rm pe}^{2}$, and the sensitivity estimation error variance $\mathbb{E}[\lVert e^{l}\rVert^{2}]$.
 Corollary 3.12 (Asymptotic convergence).
-As $|\mathcal{T}|\rightarrow\infty$, the term $K_{1}$ approaches $K_{2}$, given by
-[…]
-$4(\kappa_{2}+\kappa_{3})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}\Bigg{\}}$
-From $K_{2}$, we observe that the achievement of an asymptotically accurate solution of (5) at a local minima $(p^{*},h(p^{*},d))$ is limited by a combination of gradient estimation errors that already include opinion and clicking behaviour estimation (i.e.
-terms $\kappa_{2},\kappa_{3}$) and sensitivity estimation error, along with the variance of the dither signal $\sigma_{\rm pe}^{2}$.
-The term $\kappa_{1}$ approaches zero asymptotically.
-Moreover the upper-bound of the average of the variance of sensitivity estimation error over the trigger time instances approaches $C_{f}$ asymptotically.
-4 Numerics
+As $|\mathcal{T}|\rightarrow\infty$, the upper bound $K_{1}$ in in Theorem 3.10 approaches $K_{2}$, given by
+$K_{2}=\frac{6}{1-2\eta L^{\prime}}\Bigg{(}$ $2C_{f}(2n\gamma^{2}+M_{x}^{2}+4n^{3/2}\epsilon_{g}L_{x})+$
+$4(\kappa_{2}+\kappa_{3})+\frac{\sigma_{\rm pe}^{2}}{\eta^{2}}\Bigg{)}$
+Proof 3.13.
+Refer to Appendix 5.15.
+From $K_{2}$, we observe that the asymptotic accuracy is limited by a combination of gradient estimation errors, which include opinion and clicking behaviour estimation (i.e., the terms $\kappa_{2},\kappa_{3}$), the sensitivity estimation error, and the variance of the dither signal $\sigma_{\rm pe}^{2}$.
+Moreover, the upper-bound of the average of the variance of sensitivity estimation error over the trigger time instances approaches $C_{f}$, the asymptotic sensitivity error variance.
+4 Numerical Case Studies
 4.1 Opinion Dynamics and Clicking Behaviour
@@
 Note that the choice of these clicking behaviors make the steady-state optimization problem (5) non-convex.
-As a proof of concept, we consider a platform with $n=15$ users, and randomly generate the adjacency matrix as $A\sim\mathcal{U}[0,1]^{n\times n}$, $A\mathbf{1}_{n}\leq\mathbf{1}_{n}$ and their initial opinions as $x^{0}\sim\mathcal{U}[-1,1]^{n}$.
+As a proof of concept, we consider a platform with $n=15$ users.
+We generate the adjacency matrix $A$ with some random entries zero and other entries sampled from $\mathcal{U}[0,1]$ [^2] This ensures that the non-zero weights in the graph are not insignificant.
+such that $A\mathbf{1}_{n}\leq\mathbf{1}_{n}$.
+Thus, the resulting graph need not be strongly connected.
+We generate the initial opinions as $x^{0}\sim\mathcal{U}[-1,1]^{n}$.
 Moreover, we set $\Gamma_{d}=0.5\mathrm{diag}[|x^{0}|]$.
@@
 We set the trade-off factor $\gamma=1$ in (3), thus giving equal importance to engagement maximization and polarization mitigation.
-In the online phase, we choose $\sigma_{\text{pe}}=0.07$ in (16) for the first $N/5$ time steps to facilitate sensitivity estimation and network learning, and $\sigma_{\text{pe}}=0$ afterwards to favor accurate recommendations.
+In the online phase, we choose $\sigma_{\text{pe}}=0.07$ in (16) for the first $N/5$ time steps to facilitate sensitivity estimation [^3] Sensitivity estimation can be halted after convergence on its entries.
+and network learning, and $\sigma_{\text{pe}}=0$ afterwards to favor accurate recommendations.
 In the Kalman filter (3.2.1), we initialize ^$\hat{\ell}^{0}=0.5{\rm vec}(I_{n})$ and $\Sigma^{0}=I_{n^{2}}$.
@@
 For the polarization cost $\varphi^{\rm pol}$ in (4), we set $\epsilon_{1}=-0.5$ and $\epsilon_{2}=0.5$.
-All the simulations are conducted for $N=2.5\cdot 10^{4}$ time instances and over $50$ Monte-Carlo trials, with each simulation taking around 2 seconds.
+All the simulations are conducted for $N=2.5\cdot 10^{4}$ time instances and over $50$ Monte-Carlo trials.
 Since clicks are stochastic, the sequence of clicks varies across each trial.
@@
 In Table 2, we summarize all the methods used for comparison and their attributes.
-In particular, for the oracle (method $M_{1}$), we employ the standard projected-gradient controller (6), wherein opinions can be directly measured, and the sensitivity and users’ clicking behavior is known.
+In particular, for the oracle (method $M_{1}$), we employ the standard projected-gradient controller (6), wherein opinions can be directly measured, and the sensitivity and users’ clicking behavior are both known.
 Method $M_{2}$ uses online sensitivity estimation.
@@
 In particular, for methods that involve sensitivity estimation (i.e., $M_{2}-M_{4}$), we observe a higher fluctuation for the first $N/5$ time instances due to the effect of the dither signal $\sigma_{\text{pe}}$.
-Our proposed method offers slightly inferior performances compared to the other OFO methods.
-However, it must be taken into account that method $M_{4}$ is agnostic about the analytical form of $g(p,x)$, and is completely data-driven.
+Our proposed method offers slightly inferior performances compared to the others, which is unsurprising as it is completely data-driven.
 In Figure 4, we illustrate the relative sensitivity estimation error.
@@
 The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
-Figure 4: Illustration of sensitivity estimation error in methods $M_{2}-M_{4}$.
+Figure 4: Evolution of the sensitivity estimation error in methods $M_{2}-M_{4}$.
 The bold lines represent the mean and the shaded region are the ±$\pm 1$ standard deviation across the $50$ Monte-Carlo trials.
@@
 The solid lines represent the mean and the shaded region (±$\pm 1$ standard deviation) the range of changes across the $50$ Monte Carlo trials.
-4.4 Impact of Network Aware Recommendations
-We showcase the impact of providing network-aware recommendations compared to individual, decoupled recommendations.
+4.4 Impact of Network-aware Recommendations
+In this case study, we showcase the impact of providing network-aware recommendations compared to individual, decoupled recommendations.
 To achieve this, we compare our approach with two network-agnostic methods: (i) the recommender system described in [9], and (ii) a naive version of our OFO-based recommender design.
@@
 Conversely, if the user shows an equal likelihood of clicking on both $+1$ and $-1$ positions, the recommendations will be random.
-The latter is a network-agnostic version of our recommender system design summarized in Algorithm 1.
-Specifically, for each user $i\in[n]$, we use a local ANN for opinion estimation that is trained using local positions and acceptance ratio only, i.e., $\hat{x}_{i}=\hat{\beta}_{i}(y_{i},p_{i})$ rather than $\hat{\beta}_{i}(y_{i},p)$.
+It’s important to note that the comparison with the method in [9] is not entirely equitable in terms of methodology: their approach relies solely on extreme positions and does not target polarization reduction as an objective.
+However, we include this comparison because, to the best of our knowledge, it is the only model-free, opinion agnostic, systems-theory-based recommender design available.
+The second method considered for comparison is a network-agnostic version of our recommender system design summarized in Algorithm 1.
+We will refer to this method as naive OFO.
+Specifically, for each user $i\in[n]$, we use a local ANN for opinion estimation that is trained using local positions and acceptance ratio, i.e., $\hat{x}_{i}=\hat{\beta}_{i}(y_{i},p_{i})$ rather than $\hat{\beta}_{i}(y_{i},p)$.
 Further, we do not carry out sensitivity estimation and instead use a constant diagonal sensitivity $\hat{H}\in\mathcal{D}_{(0,0.5]}^{n}$ in (17).
@@
 We compare the performance of our recommender system with the other decoupled methods in terms of final mean CTR $J:=\sum_{k=N-T+1}^{N}\lVert c^{k}\rVert_{1}/(nT)$ and final polarization cost $\varphi^{\rm pol}(x^{N})$ in Figure 7 and the fixed point residual mapping in Figure 8.
-In Figure 7, we observe that our proposed algorithm achieves a slightly higher user engagement compared to its naive implementation and results in a lower polarization cost.
+In Figure 7, we observe that our proposed algorithm achieves a slightly higher ($\sim 1.3\%$) user engagement compared to its naive implementation and results in a lower polarization cost ($\sim 9.3\%$).
 Thus, network-aware recommendations effectively reduce polarization while maintaining high user engagement.
@@
 This result is unsurprising since polarization is a global network metric, whereas engagement maximization can be addressed individually by tailoring recommended content for each user.
-The recommender design in [10], which employs only extreme recommendations, naturally leads to higher polarization costs, as shown in Figure 7(b).
+Note that the recommender design in [9] employs only extreme recommendations, naturally leading to higher polarization costs, as shown in Figure 7(b).
 Furthermore, it results in lower user engagement compared to both our algorithm and the naive method.
 This is because some users in the network are assigned clicking behavior $C_{B}$, for which extreme recommendations do not necessarily increase engagement.
-Also, the setup in [10] does not consider polarization reduction as an objective.
 5 Conclusion
-Our aim is to improve the understanding of real-world phenomena using a simplified yet insightful model.
-We designed a recommender system that simultaneously maximizes user engagement and mitigates polarization.
-Our recommender system solely relies on clicks and does not require any prior knowledge about opinion dynamics and users’ clicking behavior.
-We provided theoretical optimality and closed-loop guarantees for the resulting recommender-social network interconnection.
-Finally, our simulations demonstrated that our recommender performs favorably against other approaches that do not leverage information about users’ interconnections.
-We provide evidence that polarization risk should be considered at the recommendation level.
-Future research directions include relaxing the smoothness assumptions on the clicking behaviour and incorporate other interest attractors towards recommendation other than confirmation bias, e.g.
-repulsion.
-Finally, we aim for our network-centered perspective to enhance the existing literature on opinion polarization caused by algorithmic systems and inspire effective mitigation strategies.
+We introduced a novel recommender system algorithm based on online feedback optimization that balances the trade-off between maximizing user engagement (CTR) and minimizing polarization.
+By integrating auxiliary estimation blocks, our approach overcomes the limitations of existing OFO designs, which require unavailable real-time data, and instead relies solely on click data.
+We established theoretical guarantees for closed-loop stability and validated our design through numerical simulations using an extended FJ model.
+Our results demonstrated the impact of network-awareness in reducing polarization while maintaining high engagement.
+Future compelling research directions include relaxing the smoothness assumptions for clicking behavior and exploring alternative interest drivers, such as repulsion, beyond confirmation bias.
+Parameter estimation for clicking behavior could also be refined by using a grey-box model instead of a purely black-box approach.
+Additionally, investigating the impact of different graph topologies on polarization and developing distributed algorithms for more efficient sensitivity estimation in large-scale networks would be valuable.
+Lastly, we hope our network-centered approach contributes to the literature on mitigating opinion polarization through algorithmic systems.
 Appendix
@@
 Given that the input and output layers are linear maps, we can directly make use of [46, Theorem 7] to state the following upper-bound $\forall i\in[n]$:
-$|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|\leq$ $3\underset{y_{i}\in\mathcal{X}_{y,i},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y,p)-\hat{\beta}_{i}(y,p)|+$
+$|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|\leq$ $3\underset{y_{i}\in\mathcal{X}_{y,i},p\in\mathcal{X}_{p}}{\mathrm{sup}}|\beta_{i}(y_{i},p)-\hat{\beta}_{i}(y_{i},p)|+$
 $2\omega_{\beta_{i}}(\gamma_{x})+|v_{0,i}|\gamma_{x},$
@@
 Further, $L^{\prime}$ and $\eta$ are the smoothness factor of $\varphi$ and the step-size of the gradient descent algorithm, respectively.
-Proof 5.14.
-The proof is given in [43, Lemma 2].
+Proof 5.15.
+The proof is given in [49, Lemma 2].
 Remark 5.16 (Composite gradient lipschitzness).